### v2

# On the limiting Horn inequalities

Abstract.
The Horn inequalities characterise the possible spectra of triples of $n$-by-$n$ Hermitian matrices $A+B=C$. We study integral inequalities that arise as limits of Horn inequalities as $n\to\infty$. These inequalities are parametrised by the points of an infinite-dimensional convex body, the asymptotic Horn system $\mathscr{H}$, which can be regarded as a topological closure of the countable set of Horn inequalities for all finite $n$.
We prove three main results. The first shows that arbitrary points of $\mathscr{H}$ can be well approximated by specific sets of finite-dimensional Horn inequalities. The second is a quantitative result on the redundancy of the Horn inequalities in an infinite-dimensional setting. Concretely, the Horn inequalities for finite $n$ are indexed by certain sets $T^{n}_{r}$ with $11\leq r\leq n-$ 1; we show that if $(n_{k})_{k\geq 1}$ and $(r_{k})_{k\geq 1}$ are any sequences such that $(r_{k}/n_{k})_{k\geq 1}$ is a dense subset of $(0,1)$, then the Horn inequalities indexed by the sets $T^{n_{k}}_{r_{k}}$ are sufficient to imply all of the others.
Our third main result shows that $\mathscr{H}$ has a remarkable self-characterisation property. Namely, membership in $\mathscr{H}$ is determined by the very inequalities corresponding to the points of $\mathscr{H}$ itself. To illuminate this phenomenon, we sketch a general theory of sets that characterise themselves in the sense that they parametrise their own membership criteria, and we consider the question of what further information would be needed in order for this self-characterisation property to determine the Horn inequalities uniquely.
Key words and phrases:
Eigenvalues, Hermitian matrices, Horn inequalities, Self-characterisation
1991 Mathematics Subject Classification:
Primary: 15A42. Secondary: 51M20, 14N15, 03B60
1. Introduction
Horn’s problem is a fundamental question in linear algebra: what are the possible spectra of three $n$-by-$n$ Hermitian matrices satisfying $A+B=C$? The answer to this question was first conjectured by Horn [20] and proved in a culmination of works by multiple authors, most notably Klyachko [21] and Knutson and Tao [23]. Briefly, given $\alpha,\beta,\gamma\in\mathbb{R}^{n}$, there is a recursive procedure for enumerating a system of linear inequalities on the triple $(\alpha,\beta,\gamma)\in\mathbb{R}^{3n}$, and these inequalities are a necessary and sufficient condition for the existence of three such matrices $A+B=C$ with $\alpha,\beta,\gamma$ as their respective spectra.
Simultaneously with the solution of the finite-dimensional Horn’s problem, various authors began to consider infinite-dimensional versions that deal with operators on a Hilbert space. Here there are a number of ways to pose the problem by placing different stipulations on the operators and their spectral data. One strand of work has dealt with the case of compact operators [16, 19, 4, 5], while many other papers have considered the setting of von Neumann algebras, specifically finite factors [2, 3, 1, 8, 9, 10]. In this latter case, one associates to each self-adjoint operator a probability measure on its spectrum (its *-distribution), and one asks what *-distributions can arise for operators $A+B=C$. The solution is described by a system of infinitely many inequalities on an infinite-dimensional space of triples of probability measures. Specifically, one can write the original Horn inequalities for $n$-by-$n$ matrices as integral inequalities on the empirical spectral measures; then the possible *-distributions for self-adjoint operators $A+B=C$ are those that simultaneously satisfy the Horn inequalities for all finite $n$ [3, 1].
In this paper, we take a different perspective on the asymptotics of Horn’s problem as $n\to\infty$. Here we sidestep questions about spectra of operators and focus instead on the inequalities themselves. Our investigation begins from three striking observations about the infinite system of Horn inequalities mentioned above.
The first observation is that there are too many inequalities. The list enumerated by Horn’s algorithm is known to be enormously redundant. This is true even in finite dimensions, but in infinite dimensions the phenomenon is more dramatic due to the fact that any given Horn inequality is implied asymptotically by approximation with sequences of other inequalities. In particular, the set of solutions does not change if any finite number of inequalities are discarded. We would like to know which subsets of inequalities are sufficient to characterise the problem and which can be safely ignored without changing the solutions. In the infinite-dimensional setting, such subsets can have a completely different character than in finite dimensions.
The second observation is that there are too few inequalities. Solutions to the Horn inequalities also solve many other inequalities that are not in the list. For example, suppose that $(\pi_{1},\pi_{2},\pi_{3})$ are a triple of compactly supported probability measures satisfying the Horn inequalities, and write $Q_{i}:[0,1]\to\mathbb{R}$ for the quantile function of $\pi_{i}$. (We recall the definition of quantile functions in Section 2.1.) Then the Horn inequalities imply the limiting Ky Fan inequalities:
(1.1) $\int_{x}^{1}Q_{1}(t)+Q_{2}(t)-Q_{3}(t)\,\mathrm{d}t\geq 0\qquad\forall\,x\in[0,1].$
However, the Horn inequalities themselves are countable and only include the cases where $x\in\mathbb{Q}\cap(0,1)$. Clearly, therefore, it is both possible and desirable to take some kind of closure or completion of the Horn inequalities in order to obtain a much larger list of criteria that the solutions must satisfy. But in precisely what sense can we take such a closure while ensuring that the resulting inequalities remain valid for all solutions to the original system? This turns out to be a more delicate issue than the simple example of (1.1) might suggest.
The third observation is that the set of all Horn inequalities describes itself in a peculiar sense. The Horn inequalities for a triple $(\alpha,\beta,\gamma)\in\mathbb{R}^{3n}$ all take the form
(1.2) $\sum_{i\in I}\alpha_{i}+\sum_{j\in J}\beta_{j}\geq\sum_{k\in K}\gamma_{k}$
for certain subsets $I,J,K\subset\{1,\ldots,n\}$ with $|I|=|J|=|K|=r$. There is a natural way to index such inequalities by vectors $(\lambda(I),\lambda(J),\lambda(K))\in\mathbb{R}^{3r}$, and it turns out that the triples of subsets $(I,J,K)$ corresponding to Horn inequalities are precisely those for which $(\lambda(I),\lambda(J),\lambda(K))$ solves the Horn inequalities. This fact is a straightforward consequence of Horn’s recursive procedure, but in infinite dimensions it takes a stranger form: each Horn inequality is indexed by a triple of probability measures that solves all of the others, not merely those for some fixed $n$. Thus the full set of inequalities encodes a set of necessary and sufficient conditions for membership in itself. It is natural to ask, does this self-characterisation property still hold for the “completion” of the Horn inequalities discussed above? And, how close does this property come to uniquely determining the Horn inequalities?
Here we gain new insights into all of the above questions by identifying the Horn inequalities with points in an infinite-dimensional space and studying their topological closure, which represents the set of integral inequalities on triples of probability measures that can be obtained as scaling limits of Horn inequalities as $n$ grows large. This set can be regarded either as a system of inequalities or as a geometric object in its own right, and we consider it from both perspectives.
Concretely, the central object of our investigation is the asymptotic Horn system $\mathscr{H}$, which consists of all triples of probability measures that can be obtained as weak limits of empirical spectral measures of $n$-by-$n$ Hermitian triples $A_{n}+B_{n}=C_{n}$ with eigenvalues in $[0,1]$. As we explain, points of $\mathscr{H}$ correspond to families of integral inequalities that arise as limits of Horn inequalities in a suitable topology. Thus $\mathscr{H}$ can be regarded as the closure of the Horn inequalities when viewed “from infinity.”
Our first main result says, intuitively, that arbitrary elements of $\mathscr{H}$ can be approximated by certain subsets of $\mathscr{H}$ corresponding to a sparse selection of the finite-dimensional Horn inequalities. As a consequence, we obtain our second main result, which gives a quantitative bound on the redundancy in the Horn inequalities as $n\to\infty$, in the following sense. Write $T^{n}_{r}$ for the set of triples $(I,J,K)\subset\{1,\ldots,n\}$ with $|I|=|J|=|K|=r$ that yield valid Horn inequalities for $n$-by-$n$ matrices as in (1.2). Let $(r_{k})_{k\geq 1}$ and $(n_{k})_{k\geq 1}$ be any sequences of positive integers such that $(r_{k}/n_{k})_{k\geq 1}$ is a dense subset of $(0,1)$. Then the inequalities indexed by the sets $T^{n_{k}}_{r_{k}}$ for $k\geq 1$ imply the full system of Horn inequalities. This result is “quantitative” because it tells us about the rate at which Horn inequalities may be discarded without affecting the solutions of the system — or, more precisely, it shows that there is no bound on the rate, in that $n_{k}$ can grow arbitrarily quickly and $r_{k}/n_{k}$ can approximate points of $(0,1)$ arbitrarily slowly.
Our third main result shows that $\mathscr{H}$ indeed has a self-characterisation property. That is, similarly to the original countable collection of Horn inequalities, their closure $\mathscr{H}$ also describes itself: the points of $\mathscr{H}$ are in correspondence with a set of necessary and sufficient conditions that determine whether a given triple of probability measures belongs to $\mathscr{H}$. Put differently, elements of $\mathscr{H}$ are solutions of the asymptotic Horn inequalities, but in another sense they also are the asymptotic Horn inequalities. This property is an asymptotic residue of the recursive structure of Horn’s problem in finite dimensions, whereby the inequalities for $n$-by-$n$ matrices can be constructed from the inequalities for $r$-by-$r$ matrices with $r\leq n-1$.
The self-characterisation property also answers the question of how the Horn inequalities can be “completed”: it implies that any triple of compactly supported probability measures that satisfies the Horn inequalities in fact satisfies all inequalities corresponding to points in $\mathscr{H}$. It is not obvious that such a property would persist after taking the topological closure: although the solutions to the Horn inequalities form a closed set, the inequalities themselves do not, and they are defined in terms of a functional that is not continuous in the relevant topology. Therefore it is a nontrivial fact that the solutions of the original system actually satisfy all of the inequalities in the closure.
At the end of the paper, we ask: what additional information, if any, would be needed in order for the self-characterisation property to uniquely determine the Horn inequalities? To illuminate this question and provide some partial answers, we sketch a general theory of sets that characterise themselves, in the sense that elements of the set parametrise a collection of necessary and sufficient conditions for membership in the set. We prove some abstract but elementary results on the existence, uniqueness, and structure of such sets in a very general setting, and we then apply these to the particular case of the Horn inequalities.
In the course of our analysis we will draw extensively on the theorems and techniques of authors such as Knutson and Tao [23] and Bercovici and Li [3]. We stress that we do not aim to reprove their results, but rather to apply them to new asymptotic questions about the Horn inequalities, which we hope will offer new insights into the finite-dimensional setting as well.
1.1. Overview
The outline of the paper is as follows.
Section 2 presents our main results and introduces definitions and notation that we will use throughout the article. In particular, we define our main object of study, the asymptotic Horn system $\mathscr{H}$, as well as a variant, the extended asymptotic Horn system $\mathscr{H}^{\prime}$, and we discuss some of their basic properties.
In Section 3 we review the Horn inequalities for $n$-by-$n$ matrices, and then reformulate them in an infinite-dimensional setting as integral inequalities on triples of quantile functions of probability measures, rather than linear inequalities on triples of vectors. This yields a uniform statement of the Horn inequalities that applies independently of the value of $n$ and is thus well-suited for studying asymptotic problems.
Section 4 contains the proofs of our main theorems on self-characterisation and approximation of elements of $\mathscr{H}$, as well as some corollaries and auxiliary results.
In Section 5 we begin to develop a general theory of self-characterising sets, and we consider the question of how close the self-characterisation property comes to determining the Horn inequalities uniquely.
Finally, in Section 6, we discuss some possible directions for further research.
2. The asymptotic Horn system
2.1. Definitions
The goal of this article is to prove some fundamental properties of the asymptotic Horn system $\mathscr{H}$. This is the set defined as follows. Given an $n$-by-$n$ Hermitian matrix $A$, its empirical spectral measure is the probability measure
$\pi_{A}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{\alpha_{i}},$
where $\alpha_{1}\leq\ldots\leq\alpha_{n}\in\mathbb{R}$ are the eigenvalues of $A$. We say a probability measure $\pi$ is $n$-atomic if it can be written $\pi=\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}}$. Let $\mathscr{H}_{n}$ be the collection of triples $(\pi_{1},\pi_{2},\pi_{3})$ of $n$-atomic probability measures $(\pi_{1},\pi_{2},\pi_{3})$ that are the respective empirical spectra of $n$-by-$n$ Hermitian matrices $A_{1},A_{2},A_{3}$ satisfying $A_{1}+A_{2}=A_{3}$. Let $\mathcal{P}_{[0,1]}$ be the set of probability measures with support contained in the closed unit interval $[0,1]$. We define
$\mathscr{H}:=\text{Closure of }\bigcup_{n\geq 1}\mathscr{H}_{n}\cap\mathcal{P}_{[0,1]}^{3}$
in the product topology given by weak convergence of probability measures on $\mathbb{R}$. That is, a triple of probability measures $(\pi_{1},\pi_{2},\pi_{3})$ lies in $\mathscr{H}$ if and only if there exists a sequence of triples $(\pi_{1,k},\pi_{2,k},\pi_{3,k})_{k\geq 1}$ with $(\pi_{1,k},\pi_{2,k},\pi_{3,k})$ in some $\mathscr{H}_{n_{k}}$, such that for $i=1,2,3$, ${\rm supp}\pi_{i,k}\subseteq[0,1]$ and $\lim_{k\to\infty}f(\pi_{i,k})=f(\pi_{i})$ for all bounded continuous $f:\mathbb{R}\to\mathbb{R}$.
We also define the extended asymptotic Horn system
$\mathscr{H}^{\prime}:=\text{Closure of }\bigcup_{n\geq 1}\mathscr{H}_{n}$
in the product topology obtained from the Wasserstein metric on probability measures with finite mean. More explicitly, recall that for two probability measures $\mu,\nu$ on $\mathbb{R}$ with finite mean, their Wasserstein distance is defined by
(2.1) $W_{1}(\mu,\nu):=\inf_{\gamma\in\Gamma(\mu,\nu)}\mathbb{E}_{(x,y)\sim\gamma}\big{[}|x-y|\big{]},$
where $\Gamma(\mu,\nu)$ is the space of couplings of $\mu$ and $\nu$. Convergence in Wasserstein distance is equivalent to weak convergence plus convergence in expectation, and is also equivalent to $L^{1}$ convergence of quantile functions (defined below). A triple of probability measures $(\pi_{1},\pi_{2},\pi_{3})$ lies in $\mathscr{H}^{\prime}$ if and only if there exists a sequence of triples $(\pi_{1,k},\pi_{2,k},\pi_{3,k})_{k\geq 1}$ with $(\pi_{1,k},\pi_{2,k},\pi_{3,k})$ in some $\mathscr{H}_{n_{k}}$, such that $\lim_{k\to\infty}W_{1}(\pi_{i,k},\pi_{i})=0$ for $i=1,2,3$.
We refer to $\mathscr{H}$ as the asymptotic Horn system for two reasons. The first reason we will discuss in detail below: the points of $\mathscr{H}$ are in correspondence with a system of integral inequalities that can be regarded as a metric completion or topological closure of the Horn inequalities. (The same is not true for all points of $\mathscr{H}^{\prime}$.) The second reason is to distinguish the objects $\mathscr{H}$ and $\mathscr{H}^{\prime}$ from the related asymptotic Horn bodies that have been studied elsewhere in the literature [8, 10], which are sets of the form
(2.2) $\mathscr{H}_{\mu,\nu}:=\big{\{}(\pi_{1},\pi_{2},\pi_{3})\in\mathscr{H}^{\prime}\ |\ \pi_{1}=\mu,\,\pi_{2}=\nu\big{\}}$
for $\mu,\nu$ two fixed probability measures.
Since weak convergence plus uniformly bounded support implies convergence in Wasserstein distance, we have $\mathscr{H}\subset\mathscr{H}^{\prime}$. Observe also that, given Hermitian matrices $A+B=C$, we have not only $(\pi_{A},\pi_{B},\pi_{C})\in\mathscr{H}_{n}$ but additionally $(\pi_{sA+tI},\pi_{sB+uI},\pi_{sC+(t+u)I})\in\mathscr{H}_{n}$ as well for all $s,t,u\in\mathbb{R}$, where $I$ is the $n$-by-$n$ identity matrix. It follows that $\mathscr{H}$’ is invariant under transformations of the form $(\pi_{1},\pi_{2},\pi_{3})\mapsto(T_{\#}^{t}D_{\#}^{s}\pi_{1},T_{\#}^{u}D_{\#}^{s}\pi_{2},T_{\#}^{t+u}D_{\#}^{s}\pi_{3})$, where $D^{s}(x)=sx$ and $T^{t}(x)=x+t$ for $x\in\mathbb{R}$, and $S_{\#}\pi$ indicates the pushforward of a measure $\pi$ under a measurable map $S:\mathbb{R}\to\mathbb{R}$, i.e. $S_{\#}\pi(E):=\pi(S^{-1}(E))$. Moreover, any compactly supported triple in $\mathscr{H}^{\prime}$ can be obtained from a triple in $\mathscr{H}$ via such a transformation. Thus it is quite natural and non-restrictive that we consider only measures supported on $[0,1]$ in the definition of $\mathscr{H}$.
We will also find occasion in the sequel to consider the subset $\mathscr{H}[0,1-\mu)$ consisting of those triples with support contained in $[0,1-\mu)$ for some $\mu\geq 0$.
Our results deal with various conditions that characterise the triples $(\pi_{1},\pi_{2},\pi_{3})$ belonging to $\mathscr{H}$ or $\mathscr{H}^{\prime}$. One can observe immediately that by the trace equality $\operatorname{tr}(A)+\operatorname{tr}(B)=\operatorname{tr}(C)$, any such triple must satisfy
(2.3) $0=-\int_{-\infty}^{\infty}x\,\pi_{1}(\mathrm{d}x)-\int_{-\infty}^{\infty}x\,\pi_{2}(\mathrm{d}x)+\int_{-\infty}^{\infty}x\,\pi_{3}(\mathrm{d}x).$
There are, of course, far more demanding requirements a triple of probability measures must meet in order to guarantee membership in $\mathscr{H}$ or $\mathscr{H}^{\prime}$. To describe these requirements, it is helpful to consider probability measures in terms of their quantile functions. Given a probability measure $\pi$ on $\mathbb{R}$, we define its quantile function $Q:[0,1]\to\mathbb{R}\cup\{\pm\infty\}$ to be the unique right-continuous nondecreasing function satisfying
$t=\int_{-\infty}^{Q(t)}\pi(\mathrm{d}x)\qquad\text{for $t\in[0,1]$}.$
If $\pi$ has a strictly increasing and continuous cumulative distribution function $F$, then $Q=F^{-1}$. Otherwise the quantile function may be discontinuous if $\pi$ has gaps in its support, and it may be constant over open intervals if $\pi$ has atoms. If $\pi$ is supported in an interval $[a,b]$, then $Q(t)$ takes values in $[a,b]$. The quantile function has the property that for bounded and measurable $f$ we have
(2.4) $\int_{-\infty}^{\infty}f(x)\,\pi(\mathrm{d}x)=\int_{0}^{1}f(Q(t))\,\mathrm{d}t.$
In other words, $\pi$ is the pushforward by $Q$ of Lebesgue measure on $[0,1]$.
Since the correspondence between probability measures and their quantile functions is bijective, we may equally think of $\mathscr{H}$ or $\mathscr{H}^{\prime}$ as consisting of triples $\mathbf{Q}=(Q_{1},Q_{2},Q_{3})$ of quantile functions. Everywhere below, the relevant notion of convergence for such triples can be understood as Wasserstein convergence of probability measures in each entry or, equivalently, $L^{1}$ convergence of quantile functions. We will freely use the identification of measures with their quantile functions in the sequel.
Given two triples of quantile functions $\mathbf{Q}=(Q_{1},Q_{2},Q_{3})$ and $\mathbf{\tilde{Q}}=(\tilde{Q}_{1},\tilde{Q}_{2},\tilde{Q}_{3})$ with each $\tilde{Q}_{i}$ taking values in $[0,1]$, define the composition functional
(2.5) $\mathcal{E}(\mathbf{Q},\mathbf{\tilde{Q}}):=\int_{0}^{1}-Q_{1}(\tilde{Q}_{1}(t))-Q_{2}(\tilde{Q}_{2}(t))+Q_{3}(\tilde{Q}_{3}(t))\,\mathrm{d}t,$
provided the integral in (2.5) is defined; if this integral is undefined, then so is $\mathcal{E}(\mathbf{Q},\mathbf{\tilde{Q}})$. The composition functional plays a fundamental role throughout this article. It is linear in its first argument, in that for any triples $\mathbf{P}$ and $\mathbf{Q}$ of quantile functions and any $a,b\in\mathbb{R}$, we have
(2.6) $\mathcal{E}(a\mathbf{Q}+b\mathbf{P},\mathbf{\tilde{Q}})=a\mathcal{E}(\mathbf{Q},\mathbf{\tilde{Q}})+b\mathcal{E}(\mathbf{P},\mathbf{\tilde{Q}}).$
On the other hand, if we regard the arguments of $\mathcal{E}$ as triples of measures rather than quantile functions, then $\mathcal{E}(\bm{\pi},\tilde{\bm{\pi}})$ is linear in its second argument in terms of affine addition of measures. Indeed, note that $\int_{0}^{1}Q_{i}(\tilde{Q}_{i}(t))\mathrm{d}t=\int_{0}^{1}Q_{i}(x)\tilde{\pi}_{i}(\mathrm{d}x)$. It follows that for any triples $\tilde{\bm{\pi}}$ and $\tilde{\bm{\sigma}}$ of probability measures supported on $[0,1]$ and any $\lambda\in[0,1]$, we have
(2.7) $\mathcal{E}(\bm{\pi},\lambda\tilde{\bm{\pi}}+(1-\lambda)\tilde{\bm{\sigma}})=\lambda\mathcal{E}(\bm{\pi},\tilde{\bm{\pi}})+(1-\lambda)\mathcal{E}(\bm{\pi},\tilde{\bm{\sigma}}).$
Letting $\mathbf{t}$ denote the triple of quantile functions $Q_{1}(t)=Q_{2}(t)=Q_{3}(t)=t$, by setting $f(x)=x$ in (2.4), the trace equality (2.3) then reads
(2.8) $\mathrm{tr}(\mathbf{Q}):=\mathcal{E}(\mathbf{Q},\mathbf{t})=\mathcal{E}(\mathbf{t},\mathbf{Q})=\int_{0}^{1}-Q_{1}(t)-Q_{2}(t)+Q_{3}(t)\,\mathrm{d}t=0.$
Given a triple of quantile functions $\mathbf{Q}=(Q_{1}(t),Q_{2}(t),Q_{3}(t))$ and $\mu\geq 0$, we can create a new triple
$\mathbf{Q}+\mu\mathbf{t}:=(Q_{1}(t)+\mu t,Q_{2}(t)+\mu t,Q_{3}(t)+\mu t).$
Finally, define
(2.9) $\eta_{\mathbf{Q}}:=1-\max_{i=1,2,3}\sup_{t\in[0,1]}Q_{i}(t).$
If the triple of probability measures corresponding to $\mathbf{Q}$ are supported on $[0,1]$, then so are the measures corresponding to $\mathbf{Q}+\mu\mathbf{t}$ for $0\leq\mu\leq\eta_{\mathbf{Q}}$. Note that $\mathrm{tr}(\mathbf{Q}+\mu\mathbf{t})=\mathrm{tr}(\mathbf{Q})-\mu/2\neq\mathrm{tr}(\mathbf{Q})$ if $\mu\neq 0$.
2.2. Main results
With the above definitions in hand, we can state our main results on approximation in $\mathscr{H}$ and self-characterisation.
Our first main result addresses the question of how well a particular subset of finite-dimensional Horn inequalities can approximate an arbitrary point of $\mathscr{H}$. As we recall in more detail in Section 3, the Horn inequalities for $n$-by-$n$ matrices are in correspondence with certain triples of subsets $I,J,K\subset\{1,\ldots,n\}$ with $|I|=|J|=|K|=r<n$. We write $T^{n}_{r}$ for the set of such triples corresponding to Horn inequalities. Below, we will define a mapping that identifies a triple $(I,J,K)\in T^{n}_{r}$ with a triple $(Q_{I,n},Q_{J,n},Q_{K,n})$ of quantile functions in $\mathscr{H}$. The theorem says that, under this identification, only Horn inequalities with a fixed asymptotic ratio $r_{n}/n\to q$ are needed to approximate any point in $\mathscr{H}$.
Theorem 2.1.
Let $\mathbf{Q}\in\mathscr{H}$. Then for every $0\leq q\leq\eta_{\mathbf{Q}}$ there exists a sequence $((I_{n},J_{n},K_{n}))_{n\geq 1}$ such that $(I_{n},J_{n},K_{n})\in T^{n}_{r_{n}}$ with $r_{n}/n\to q$ and $(Q_{I_{n},n},Q_{J_{n},n},Q_{K_{n},n})\to\mathbf{Q}$.
In fact we infer Theorem 2.1 from a more detailed result, which is one of the key technical tools in the paper (Theorem 4.4). It gives a quantitative bound on the distance between points of $\mathscr{H}$ and triples representing Horn inequalities in specific sets $T^{n}_{r}$.
Our second main result concerns the redundancy of the Horn inequalities as $n\to\infty$. The full statement is given below in Theorem 4.7. To keep the notation light, for now we give only an informal statement.
Theorem 2.2 (Informal version of Theorem 4.7).
Fix any sequences $(n_{k})_{k\geq 1}$, $(r_{k})_{k\geq 1}$ of positive integers such that $(r_{k}/n_{k})_{k\geq 1}$ is a dense subset of $(0,1)$. Then, in the quantile function formulation of the Horn inequalities developed in Section 3 (which applies uniformly for all values of $n$), it is sufficient to consider only Horn inequalities indexed by the sets $T^{n_{k}}_{r_{k}}$. All of the other inequalities are redundant and are implied by this subset.
Our third main result is the theorem on self-characterisation of $\mathscr{H}$.
Theorem 2.3 (The self-characterisation property).
Let $\mathbf{Q}=(Q_{1},Q_{2},Q_{3})$ be a triple of quantile functions in $L^{1}([0,1])$ satisfying $\mathrm{tr}(\mathbf{Q})=0$. Then
$\mathbf{Q}\in\mathscr{H}^{\prime}$ $\iff\mathcal{E}(\mathbf{Q},\mathbf{\tilde{Q}}+\mu\mathbf{t})\geq 0\text{ for every $\mathbf{\tilde{Q}}\in\mathscr{H}[0,1)$ and $0<\mu\leq\eta_{\mathbf{\tilde{Q}}}$.}$
Moreover, if each $Q_{i}$ takes values in $[0,1]$, then
$\mathbf{Q}\in\mathscr{H}$ $\iff\mathcal{E}(\mathbf{Q},\mathbf{\tilde{Q}}+\mu\mathbf{t})\geq 0\text{ for every $\mathbf{\tilde{Q}}\in\mathscr{H}$ and $0\leq\mu\leq\eta_{\mathbf{\tilde{Q}}}$.}$
This theorem says that each point in $\mathscr{H}$ (resp. $\mathscr{H}[0,1)$) determines a family of integral inequalities, and that all of these inequalities taken together are a necessary and sufficient condition for a suitable triple of quantile functions to belong to $\mathscr{H}$ (resp. $\mathscr{H}^{\prime}$). This property arises in the $n\to\infty$ limit from the recursive structure of the Horn inequalities, and it means, in a specific sense, that the set $\mathscr{H}$ encodes a precise description of itself. It also confirms that the compactly supported solutions of the full asymptotic Horn system are the same as those of the countable family of finite-$n$ Horn inequalities. This provides an answer to the question of how the Horn inequalities can be “completed.”
Remark 2.4.
As we explain below, if $\mathbf{Q}$ is a triple of quantile functions of empirical spectral measures of $n$-by-$n$ matrices, and if $\mathbf{\tilde{Q}}=(Q_{I,n},Q_{J,n},Q_{K,n})$ for some triple $(I,J,K)\in T^{n}_{r}$, then then corresponding Horn inequality can be expressed as
$\mathcal{E}\Big{(}\mathbf{Q},\mathbf{\tilde{Q}}+\frac{r}{n}\mathbf{t}\Big{)}\geq 0.$
Thus the parameter $\mu$ appearing in the inequalities in Theorem 2.3 should be thought of as a limiting value of the ratio $r/n$.
Remark 2.5.
For particular choices of triples, Theorem 2.3 recovers various classical inequalities as well as their asymptotic analogues. We have already seen one example above in the limiting Ky Fan inequalities (1.1). For another example, observe that for any $a,b\in\mathbb{R}$, the triple $(\delta_{a},\delta_{b},\delta_{a+b})$ of Dirac masses lies in $\mathscr{H}^{\prime}$, and their respective quantile functions are $\tilde{Q}_{1}(t)=a,\tilde{Q}_{2}(t)=b,\tilde{Q}_{3}(t)=a+b$. The corresponding inequalities tell us that for any $a,b$ with $0\leq a,b,a+b\leq 1$, each triple $\mathbf{Q}=(Q_{1}(t),Q_{2}(t),Q_{3}(t))\in\mathscr{H}^{\prime}$ must satisfy
(2.10) $-Q_{1}(a)-Q_{2}(b)+Q_{3}(a+b)\geq 0.$
This is an asymptotic analogue of the Weyl inequalities [30] and was previously observed by Bercovici and Li, who also used large-$n$ limits of Horn inequalities to derive an infinite-dimensional version of the Freede–Thompson inequalities in a similar fashion [2].
Alternatively, let $\mu\in[0,1]$ and fix any quantile function $S:[0,1]\to[0,1-\mu]$, and consider the element $(\tilde{Q}_{1}(t),\tilde{Q}_{2}(t),\tilde{Q}_{3}(t))=(S(t),0,S(t))\in\mathscr{H}$. Then any triple $\mathbf{Q}=(Q_{1}(t),Q_{2}(t),Q_{3}(t))$ must satisfy
(2.11) $-\int_{0}^{1}Q_{1}(S(t)+\mu t)\,\mathrm{d}t-\int_{0}^{1}Q_{2}(\mu t)\,\mathrm{d}t+\int_{0}^{1}Q_{3}(S(t)+\mu t)\,\mathrm{d}t\geq 0.$
This is an asymptotic analogue of the Lidskii–Wielandt inequality [25, 31].
Naturally, it is possible to obtain asymptotic analogues of various other inequalities by considering different limiting Horn triples.
Remark 2.6.
Theorem 2.3 in fact implies that whenever $\mathbf{Q}\in\mathscr{H}^{\prime}$ represents a triple of compactly supported probability measures, we have $\mathcal{E}(\mathbf{Q},\mathbf{\tilde{Q}}+\mu\mathbf{t})\geq 0$ for every $\mathbf{\tilde{Q}}\in\mathscr{H}$ and $0\leq\mu\leq\eta_{\mathbf{\tilde{Q}}}$ (see Corollary 4.9 below). On the other hand, if some measure represented by $\mathbf{Q}$ has unbounded support, then the corresponding quantile function is also unbounded in a neighborhood of 0 and/or 1. In that case, for $\mathbf{\tilde{Q}}\in\mathscr{H}$ corresponding to a triple of probability measures that do not have bounded densities near the endpoints of the unit interval, the integrals in the definition (2.5) may diverge, and thus $\mathcal{E}(\mathbf{Q},\mathbf{\tilde{Q}})$ may not be defined. The requirement that $\mathbf{\tilde{Q}}\in\mathscr{H}[0,1)$ and $0<\mu\leq\eta_{\mathbf{\tilde{Q}}}$ guarantees that $\mathbf{\tilde{Q}}+\mu\mathbf{t}$ represents a triple of measures with densities bounded above by $1/\mu$, which in turn guarantees that $\mathcal{E}(\mathbf{Q},\mathbf{\tilde{Q}}+\mu\mathbf{t})$ is finite. This is the only reason that different sets of inequalities appear in the criteria for membership in $\mathscr{H}$ and $\mathscr{H}^{\prime}$ in Theorem 2.3.
Remark 2.7.
Readers familiar with the work of Bercovici and Li [3] may be aware that they gave an $n$-independent formulation of the Horn inequalities by identifying the collection of inequalities for all finite $n$ with a set $\mathcal{T}$ of triples of indicator functions of subsets of $[0,1]$, and that the inequalities parametrised by $\mathcal{T}$ are also solved by the elements of $\mathcal{T}$ itself — a property of a similar flavour to Theorem 2.3. Although our scheme for embedding the Horn inequalities in an infinite-dimensional space is different from that used in [3], nevertheless, if one normalises the indicator functions in [3] so that they become densities of probability measures, then the asymptotic Horn system $\mathscr{H}$ is the closure of $\mathcal{T}$ in the topology of weak convergence. Accordingly one might wonder whether Theorem 2.3 could be established via the constructions of [3] merely by proving that the functional $\mathcal{E}$ is continuous in some appropriate sense. We emphasise that this is not the case; in fact $\mathcal{E}$ is not continuous in our chosen topology, and even if it were, that would not immediately prove the desired result without further knowledge of the set $\mathscr{H}$. Although the continuity properties of $\mathcal{E}$ are a crucial ingredient in the proof of Theorem 2.3, a more delicate density argument at the level of the Horn inequalities themselves is also required.
In the course of proving the results above, we also prove the following “self-averaging” property of $\mathscr{H}$. It is not difficult to show, and an equivalent fact was already remarked in passing in [3], but we believe that it is interesting enough to merit a distinct statement and proof.
Proposition 2.8 (The self-averaging property).
Let $\mathbf{Q}=(Q_{1}(t),Q_{2}(t),Q_{3}(t))$ be a triple of quantile functions in $L^{1}([0,1])$ with $\operatorname{tr}(\mathbf{Q})=0$, such that $\mathcal{E}(\mathbf{Q},\mathbf{\tilde{Q}}+\mu\mathbf{t})\geq 0$ for all $\mathbf{\tilde{Q}}\in\mathscr{H}[0,1)$ and $0<\mu\leq\eta_{\mathbf{\tilde{Q}}}$. Define $\mathbf{Q}^{n}=(Q^{n}_{1}(t),Q^{n}_{2}(t),Q^{n}_{3}(t))$ by averaging $\mathbf{Q}$ over each interval of the form $\left[\frac{j-1}{n},\frac{j}{n}\right)$, that is,
(2.12) $Q_{i}^{n}(t):=n\int_{(j-1)/n}^{j/n}Q_{i}(u)\,\mathrm{d}u\qquad\textrm{for }t\in\left[\frac{j-1}{n},\frac{j}{n}\right),\quad 1\leq j\leq n.$
Then there exist $n$-by-$n$ Hermitian matrices $A+B=C$ with spectral quantiles given by $\mathbf{Q}^{n}$.
Note that by Theorem 2.3, the conditions on $\mathbf{Q}$ in Proposition 2.8 are equivalent to assuming $\mathbf{Q}\in\mathscr{H}^{\prime}$. Thus for each $n$, the averaging map is a linear projection on $L^{1}([0,1])$, and the above proposition says that it gives a surjection $\mathscr{H}^{\prime}\twoheadrightarrow\mathscr{H}_{n}$. In particular, for each $m$ it gives a linear transformation $\mathbb{R}^{3m}\to\mathbb{R}^{3n}$ that maps spectra of $m$-by-$m$ Hermitian triples to spectra of $n$-by-$n$ Hermitian triples. Note that averaging maps for different values of $n$ do not generally commute.
Aside from the above results, we prove a number of other properties of $\mathscr{H}$ and $\mathscr{H}^{\prime}$ and take steps towards developing a general theory of sets that characterise themselves under a relation, which we use to consider under what further conditions the self-characterisation property might uniquely determine $\mathscr{H}$.
2.3. First properties
We now begin studying the asymptotic Horn system in earnest, beginning with some elementary properties.
2.3.1. Dirac masses
By adding multiples of the $n$-by-$n$ identity matrix, we see that $(\delta_{a},\delta_{b},\delta_{a+b})\in\mathscr{H}^{\prime}$ for any $a,b\in\mathbb{R}$.
2.3.2. Dilations and translations
As noted above, by multiplying a Hermitian matrix equation by a scalar, we see that if $(\pi_{1},\pi_{2},\pi_{3})$ is a triple in $\mathscr{H}$ (resp. $\mathscr{H}^{\prime}$), then the pushforward $(D^{s}_{\#}\pi_{1},D^{s}_{\#}\pi_{2},D^{s}_{\#}\pi_{3})$ of these measures under a dilation $D^{s}(x)=sx$ for $s\in[0,1]$ (resp. $s\in\mathbb{R}$) is another element of $\mathscr{H}$ (resp. $\mathscr{H}^{\prime}$). Furthermore, by adding multiples of the identity to a Hermitian matrix equation, we see that writing $T^{a}:\mathbb{R}\to\mathbb{R}$ for translation $T^{a}(x):=x+a$, for any $a,b\in\mathbb{R}$ and $(\pi_{1},\pi_{2},\pi_{3})\in\mathscr{H}^{\prime}$ we have $(T^{a}_{\#}\pi_{1},T^{b}_{\#}\pi_{2},T^{a+b}_{\#}\pi_{3})\in\mathscr{H}^{\prime}$ also.
2.3.3. Exchange of coordinates
If $(\pi_{1},\pi_{2},\pi_{3})$ is an element of $\mathscr{H}$ or $\mathscr{H}^{\prime}$, so is $(\pi_{2},\pi_{1},\pi_{3})$. In either case, $(\pi_{1},D^{-1}_{\#}\pi_{3},D^{-1}_{\#}\pi_{2})\in\mathscr{H}^{\prime}$.
The asymptotic Horn system has two different convexity properties that it inherits from distinct linear structures on two different spaces of measures on the line, as illustrated in (2.6) and (2.7): one arises from the simple addition of finite signed measures, and the other arises from addition of quantile functions of probability measures.
2.3.4. Vertical convexity
The first, “vertical” convexity property is that whenever $(\pi_{1},\pi_{2},\pi_{3})$ and $(\tilde{\pi}_{1},\tilde{\pi}_{2},\tilde{\pi}_{3})$ are any two triples of measures in $\mathscr{H}$ (resp. $\mathscr{H}^{\prime}$), for all $\lambda\in[0,1]$ we have
(2.13) $\big{(}\lambda\pi_{1}+(1-\lambda)\tilde{\pi}_{1},\,\lambda\pi_{2}+(1-\lambda)\tilde{\pi}_{2},\,\lambda\pi_{3}+(1-\lambda)\tilde{\pi}_{3}\big{)}\in\mathscr{H}\textrm{ (resp. $\mathscr{H}^{\prime}$)}.$
Here $\lambda\pi+(1-\lambda)\tilde{\pi}$ is simply a linear combination of measures, i.e. $(\lambda\pi+(1-\lambda)\tilde{\pi})(A):=\lambda\pi(A)+(1-\lambda)\tilde{\pi}(A)$ for measurable $A$.
The property (2.13) may be established by considering block diagonal matrices consisting of a $\lfloor\lambda n\rfloor$-by-$\lfloor\lambda n\rfloor$ block and a $\lceil(1-\lambda)n\rceil$-by-$\lceil(1-\lambda)n\rceil$ block along the diagonal.
In the special case where the measure $\pi_{i}$ has support lying to the left of $\tilde{\pi}_{i}$ (i.e. $\pi_{i}$ is supported on $(-\infty,x]$ and $\tilde{\pi_{i}}$ is supported on $[x,\infty)$ for some $x\in\mathbb{R}$), a convex combination of measures corresponds to a concatenation of quantile functions. That is, the quantile function of $\lambda\pi_{i}+(1-\lambda)\tilde{\pi}_{i}$ is given by $Q_{i}(t/\lambda)$ for $t<\lambda$ and $\tilde{Q}_{i}((t-\lambda)/(1-\lambda))$ for $t\geq\lambda$.
2.3.5. Horizontal convexity
The second, “horizontal” convexity property is best stated in terms of quantile functions. Namely, whenever $\mathbf{Q}=(Q_{1},Q_{2},Q_{3})$ and $\mathbf{\tilde{Q}}=(\tilde{Q}_{1},\tilde{Q}_{2},\tilde{Q}_{3})$ are (the quantile functions of) elements of $\mathscr{H}$, for $\lambda\in[0,1]$ we have
(2.14) $\lambda\mathbf{Q}+(1-\lambda)\mathbf{\tilde{Q}}\in\mathscr{H},$
while for any $\mathbf{Q},\mathbf{\tilde{Q}}\in\mathscr{H}^{\prime}$, and $a,b\geq 0$ we have
(2.15) $a\mathbf{Q}+b\mathbf{\tilde{Q}}\in\mathscr{H}^{\prime}.$
The horizontal convexity property is less obvious, but it follows from the fact that the set of triples $(\alpha,\beta,\gamma)$ of spectra of $n$-by-$n$ Hermitian matrices $A+B=C$ is a convex polyhedral cone, together with the fact that at the level of empirical measures, addition of vectors in $\mathbb{R}^{n}$ with nonincreasing coordinates corresponds to addition of quantile functions.
2.3.6. The Sudoku property
Consider the following matrix of probability measures
$\pi_{1,1}$ $\pi_{1,2}$ $\pi_{1,3}$
$\pi_{2,1}$ $\pi_{2,2}$ $\pi_{2,3}$
$\pi_{3,1}$ $\pi_{3,2}$ $\pi_{3,3}$.
The Sudoku property states that if the first two rows and all three columns of this matrix represent elements of $\mathscr{H}$, then so does the bottom row. This follows from considering matrix equations obtained by adding rows and columns of the $2$-by-$2$ matrix of matrices $(A_{n,i,j})_{n\geq 1,1\leq i,j\leq 2}$.
The extended asymptotic Horn system $\mathscr{H}^{\prime}$ includes triples of measures that play significant roles in both classical and free probability:
2.3.7. Classical probability
We say that a probability measure $\Pi$ on $\mathbb{R}^{2}$ is a coupling of $\pi_{1}$ and $\pi_{2}$ if for all Borel subsets $A$ of $\mathbb{R}$ we have $\Pi(A\times\mathbb{R})=\pi_{1}(A)$ and $\Pi(\mathbb{R}\times A)=\pi_{2}(A)$. For any compactly supported probability measures $\pi_{1}$ and $\pi_{2}$,
$(\pi_{1},\pi_{2},\pi_{3})\in\mathscr{H}^{\prime}\text{ whenever $\pi_{3}$ is the law of $X+Y$ under a coupling $\Pi$ of $\pi_{1}$ and $\pi_{2}$}.$
This property follows from considering diagonal matrices $(B_{n,i})_{n\geq 1}$ with decreasing entries and empirical spectra converging to $\pi_{i}$ ($i=1,2$), and then setting $A_{n,1}=B_{n,1}$ and $A_{n,2}=\Sigma^{n}B_{2}^{n}(\Sigma^{n})^{-1}$ for a suitable sequence of permutation matrices $(\Sigma^{n})$.
2.3.8. Free probability
Suppose that $(B_{n,i})_{n\geq 1}$ are sequences of diagonal matrices with decreasing entries and empirical spectra converging to $\pi_{i}$ ($i=1,2$). For $i=1,2$, let $(U_{n,i})_{n\geq 1}$ be independent sequences of unitary random matrices distributed according to Haar measure, and define $A_{n,i}:=U_{n,i}B_{n,i}U_{n,i}^{-1}$. Then thanks to the results of free probability (see e.g. [27]), the empirical spectrum of the random matrix sum $A_{n,1}+A_{n,2}$ converges almost surely to a probability measure $\pi_{1}\boxplus\pi_{2}$ known as the additive free convolution of $\pi_{1}$ and $\pi_{2}$. In particular, for any compactly supported $\pi_{1}$ and $\pi_{2}$ we have
$(\pi_{1},\pi_{2},\pi_{1}\boxplus\pi_{2})\in\mathscr{H}^{\prime}.$
3. The quantile function approach to the Horn inequalities
In this section, we first review the Horn inequalities for $n$-by-$n$ matrices. We then give an $n$-independent formulation of the inequalities in terms of quantile functions of probability measures, which points to natural asymptotic questions about how the Horn inequalities and their solutions behave as $n$ grows large. Such questions may be understood in terms of the geometry and topology of the asymptotic Horn system $\mathscr{H}$, and they motivate our study of this object.
As mentioned above in Section 2.2, Bercovici and Li [3] previously gave a different $n$-independent formulation of the Horn inequalities, which is closely related to the one that we present here. Their formulation was used in [3, 1] to solve the Horn problem in the infinite-dimensional setting of an arbitrary finite factor. The key difference between the approaches taken in [3] and in the present paper is that here we identify the finite-$n$ Horn inequalities with triples of probability measures rather than triples of subsets of $[0,1]$. This allows us to give a more symmetrical statement of the Horn inequalities in terms of the functional $\mathcal{E}$, whose arguments are two triples of probability measures (via their quantile functions). It then becomes possible to study the limiting behavior of the Horn inequalities themselves by considering weak or Wasserstein convergence of these measures.
3.1. The Horn inequalities for $n$-by-$n$ matrices
In their original formulation, the Horn inequalities are a system of linear inequalities on three vectors $\alpha,\beta,\gamma\in\mathbb{R}^{n}$, of the form
(3.1) $\sum_{i\in I}\alpha_{i}+\sum_{j\in J}\beta_{j}\geq\sum_{k\in K}\gamma_{k},$
where $I,J,K$ are subsets of $\{1,\ldots,n\}$ with $|I|=|J|=|K|$.
In his landmark paper [20], Horn described an algorithm for enumerating certain triples of sets $(I,J,K)$ and conjectured that the corresponding inequalities (3.1) characterised the possible spectra of sums of two Hermitian matrices with prescribed eigenvalues.
Horn constructed these triples $(I,J,K)$ inductively as follows. We index elements of such subsets in increasing order, writing e.g. $I=\{i_{1}<\ldots<i_{r}\}$. Now first define, for $n\geq 2$ and $1\leq r<n$,
(3.2) $U^{n}_{r}=\left\{(I,J,K)\ \bigg{|}\ I,J,K\subset\{1,\ldots,n\},\ |I|=|J|=|K|=r,\ \sum_{i\in I}i+\sum_{j\in J}j=\sum_{k\in K}k+\frac{r(r+1)}{2}\right\}.$
The Horn inequalities correspond to certain subsets $T^{n}_{r}\subset U^{n}_{r}$. These are defined by setting $T^{n}_{1}=U^{n}_{1}$, and for $r>1$,
(3.3) $T^{n}_{r}=\left\{(I,J,K)\in U^{n}_{r}\ \bigg{|}\ \textrm{for all }p<r\textrm{ and all }(F,G,H)\in T^{r}_{p},\ \sum_{f\in F}i_{f}+\sum_{g\in G}j_{g}\leq\sum_{h\in H}k_{h}+\frac{p(p+1)}{2}\right\}.$
Horn made the following conjecture, now a theorem; for details of the proof, we refer the reader to [21, 23, 24] and the excellent review [18].
Theorem 3.1.
For any $\alpha,\beta,\gamma\in\mathbb{R}^{n}$ with their coordinates listed in nonincreasing order, the following are equivalent:
  1. (1)
    There exist $n$-by-$n$ Hermitian matrices $A$ and $B$ such that $(A,B,A+B)$ have respective spectra $(\alpha,\beta,\gamma)$.
  2. (2)
    $\sum_{i=1}^{n}\alpha_{i}+\sum_{j=1}^{n}\beta_{j}=\sum_{k=1}^{n}\gamma_{k}$, and the inequalities (3.1) hold for all triples $(I,J,K)\in T^{n}_{r}$ for all $11\leq r\leq n-$ 1.
The definition of the sets $T^{n}_{r}$ can also be stated in terms of integer partitions $\lambda=(\lambda_{1}\geq\ldots\geq\lambda_{r})$. For any $I\subset\{1,\ldots,n\}$ with $|I|=r$, we define a corresponding partition $\lambda(I)$ with $r$ parts by
(3.4) $\lambda(I)=\big{(}i_{r}-r,i_{r-1}-(r-1),\ldots,i_{1}-1\big{)}.$
Theorem 3.1 then yields the following alternative characterisation of the sets $T^{n}_{r}$, which makes the recursive structure of the Horn inequalities more apparent (see [18, Thm. 2]).
Theorem 3.2.
Let $I,J,K\subset\{1,\ldots,n\}$ with $|I|=|J|=|K|=r$. Then $(I,J,K)\in T^{n}_{r}$ if and only if there exist $r$-by-$r$ Hermitian matrices $(A,B,A+B)$ with spectra $(\lambda(I),\lambda(J),\lambda(K))$.
3.2. Quantile functions of Horn triples
We now define a mapping from triples of sets $(I,J,K)$ in $U_{r}^{n}$ or $T_{r}^{n}$ to triples of quantile functions $(Q_{I,n},Q_{J,n},Q_{K,n})$ of probability measures on $[0,1]$. This mapping enables a uniform formulation of the Horn inequalities that applies for all $n$.
Let $I$ be a subset of $\{1,\ldots,n\}$ with $|I|=r$, $11\leq r\leq n-$ 1, and define $\lambda(I)$ as in (3.4). The empirical measure of $\frac{1}{n}\lambda(I)$ is
(3.5) $\pi_{I,n}(\mathrm{d}x):=\frac{1}{r}\sum_{s=1}^{r}\delta_{(i_{s}-s)/n}(\mathrm{d}x).$
Note that $\pi_{I,n}$ is supported on $\{0/n,\ldots,(n-r)/n\}$. The quantile function of $\pi_{I,n}$ is
(3.6) $Q_{I,n}(t)=\frac{i_{s}-s}{n},\qquad t\in\left[\frac{s-1}{r},\frac{s}{r}\right),\quad 1\leq s\leq r.$
Given a triple of sets $(I,J,K)$ we write
(3.7) $\mathbf{Q}_{I,J,K,n}:=(Q_{I,n}(t),Q_{J,n}(t),Q_{K,n}(t)).$
In the next few lemmas, we will reformulate the definitions of the sets $U_{r}^{n}$ and $T_{r}^{n}$ in terms of the associated quantile functions.
Lemma 3.3.
Let $(I,J,K)$ be a triple of subsets of $\{1,\ldots,n\}$ of cardinality $r$. Then $(I,J,K)$ lie in $U_{r}^{n}$ if and only if $\mathrm{tr}(\mathbf{Q}_{I,J,K,n})=0$.
Proof.
A brief calculation tells us that
$\int_{0}^{1}Q_{I,n}(t)\,\mathrm{d}t=\frac{1}{r}\sum_{s=1}^{r}\frac{i_{s}-s}{n}=\frac{1}{nr}\left(\sum_{i\in I}i-\frac{r(r+1)}{2}\right).$
Using the definition (2.8) of $\mathrm{tr}(\mathbf{Q}_{I,J,K,n})$ we have
$\mathrm{tr}(\mathbf{Q}_{I,J,K,n})=\frac{1}{nr}\left(\frac{r(r+1)}{2}+\sum_{k\in K}k-\sum_{i\in I}i-\sum_{j\in J}j\right),$
which is zero if and only if $\sum_{i\in I}i+\sum_{j\in J}j=\sum_{k\in K}k+\frac{r(r+1)}{2}$, which is precisely the condition for membership in $U_{r}^{n}$. ∎
Clearly every triple $(I,J,K)$ in $U_{r}^{n}$ gives rise to a triple of quantile functions with trace zero, constant on intervals of length $1/r$, and taking values in the multiples of $1/n$. In fact, this correspondence works both ways:
Lemma 3.4.
Let $\mathbf{Q}=(Q_{1}(t),Q_{2}(t),Q_{3}(t))$ be a triple of quantile functions satisfying $\mathrm{tr}(\mathbf{Q})=0$, such that each $Q_{i}(t)$ is constant for $t\in\left[\frac{j-1}{r},\frac{j}{r}\right)$, and such that for all $t\in[0,1]$ and all $i=1,2,3$, $Q_{i}(t)=m/n$ for some $0\leq m\leq n-r$. Then there exists a triple $(I,J,K)\in U_{r}^{n}$ such that $\mathbf{Q}=\mathbf{Q}_{I,J,K,n}$.
Proof.
For $1\leq s\leq r$ write $Q_{1}(t):=\frac{m_{s}}{n}$ for $t\in\left[\frac{s-1}{r},\frac{s}{r}\right)$. Now define $i_{s}:=m_{s}+s$. Then $1\leq i_{1}<i_{2}<\ldots<i_{r}\leq n$. Thus $I:=\{i_{1},\ldots,i_{r}\}$ is a subset of $\{1,\ldots,n\}$ of cardinality $r$. Note by (3.6) that $Q_{1}(t)=Q_{I,n}(t)$. Constructing sets $J$ and $K$ the same way from $Q_{2}(t)$ and $Q_{3}(t)$ respectively, we obtain a triple $(I,J,K)$ of subsets of $\{1,\ldots,n\}$ of cardinality $r$. Thus $\mathbf{Q}=\mathbf{Q}_{I,J,K,n}$, and by Lemma 3.3 the fact that $(I,J,K)$ belongs to $U_{r}^{n}$ is equivalent to the fact that $\mathrm{tr}(\mathbf{Q})=0$. ∎
Our next lemma characterises membership of $T_{r}^{n}$ in terms of quantile functions.
Lemma 3.5.
Let $(I,J,K)$ be a triple of subsets of $\{1,\ldots,n\}$ of cardinality $r$. Then $(I,J,K)$ lie in $T_{r}^{n}$ if and only if $\mathrm{tr}(\mathbf{Q}_{I,J,K,n})=0$ and
$\mathcal{E}\Big{(}\mathbf{Q}_{I,J,K,n},\mathbf{Q}_{F,G,H,r}+\frac{p}{r}\mathbf{t}\Big{)}\geq 0$
for all $(F,G,H)$ in $T_{p}^{r}$ for all $11\leq p\leq r-$ 1.
Before proving Lemma 3.5, we have the following calculation, which describes integration against composition with the quantile function $Q_{I,n}(t)+\frac{r}{n}t$.
Lemma 3.6.
Let $P$ be a quantile function in $L^{1}([0,1])$. Then
$\int_{0}^{1}P\left(Q_{I,n}(t)+\frac{r}{n}t\right)\mathrm{d}t=\frac{n}{r}\sum_{i\in I}\int_{(i-1)/n}^{i/n}P(t)\,\mathrm{d}t.$
Proof.
Write $I=\{i_{1}<\ldots<i_{r}\}$. Using (3.6) to obtain the second equality below, we have
$\int_{0}^{1}P\left(Q_{I,n}(t)+\frac{r}{n}t\right)\mathrm{d}t$ $=\sum_{s=1}^{r}\int_{(s-1)/r}^{s/r}P\left(Q_{I,n}(t)+\frac{r}{n}t\right)\mathrm{d}t$
$=\sum_{s=1}^{r}\int_{(s-1)/r}^{s/r}P\left(\frac{i_{s}-s}{n}+\frac{r}{n}t\right)\mathrm{d}t=\frac{n}{r}\sum_{s=1}^{r}\int_{(i_{s}-1)/n}^{i_{s}/n}P(t)\,\mathrm{d}t,$
as required. ∎
We now complete the proof of Lemma 3.5.
Proof of Lemma 3.5.
We compose $Q_{I,n}(t)$ with $Q_{F,r}(t)+\frac{p}{r}t$. Using Lemma 3.6 to obtain the first equality below, and (3.6) to obtain the second, we have
(3.8) $\int_{0}^{1}Q_{I,n}\left(Q_{F,r}(t)+\frac{p}{r}t\right)\mathrm{d}t=\frac{r}{p}\sum_{f\in F}\int_{(f-1)/r}^{f/r}Q_{I,n}(t)\,\mathrm{d}t=\frac{1}{p}\sum_{f\in F}\frac{i_{f}-f}{n}.$
Then using (3.8) to obtain the first equality below, and the fact that $(F,G,H)\in U_{p}^{r}$ entails $\sum_{f\in F}f+\sum_{g\in G}g=\sum_{h\in H}h+\frac{p(p+1)}{2}$ to obtain the second, we have
$\mathcal{E}\Big{(}\mathbf{Q}_{I,J,K,n},\mathbf{Q}_{F,G,H,r}+\frac{p}{r}\mathbf{t}\Big{)}$ $=\frac{1}{pn}\left(-\sum_{f\in F}(i_{f}-f)-\sum_{g\in G}(j_{g}-g)+\sum_{h\in H}(k_{h}-h)\right)$
(3.9) $=\frac{1}{pn}\left(-\sum_{f\in F}i_{f}-\sum_{g\in G}j_{g}+\sum_{h\in H}k_{h}+p(p+1)/2\right).$
Now by the definition of $T_{r}^{n}$, $(I,J,K)$ belongs to $T_{r}^{n}$ if and only if $-\sum_{f\in F}i_{f}-\sum_{g\in G}j_{g}+\sum_{h\in H}k_{h}+p(p+1)/2\geq 0$ for all relevant $(F,G,H)$. That completes the proof. ∎
3.3. The Horn inequalities in terms of quantile functions
In fact, Lemma 3.5 above indicates how we can translate all of the Horn inequalities into integral inequalities on quantile functions. Recall from (3.1) and Theorem 3.1 that elements $\alpha,\beta,\gamma$ of $\mathbb{R}^{n}$ with coordinates in nonincreasing order are the eigenvalues of $n$-by-$n$ Hermitian matrices satisfying $A+B=C$ if and only if for every triple $(I,J,K)$ in $T_{r}^{n}$ we have
$\sum_{k\in K}\gamma_{k}\leq\sum_{i\in I}\alpha_{i}+\sum_{j\in J}\beta_{j}.$
The following proposition states this condition in terms of the quantile functions of the empirical spectral measures:
Proposition 3.7.
Let $\mathbf{Q}=(Q_{1}(t),Q_{2}(t),Q_{3}(t))$ be a triple of quantile functions that are constant on the half-open intervals $\left[\frac{j-1}{n},\frac{j}{n}\right)$, $1\leq j\leq n$. Then $\mathbf{Q}$ are the quantile functions of the empirical spectra of $n$-by-$n$ Hermitian matrices $A+B=C$ if and only if $\mathcal{E}(\mathbf{Q},\mathbf{Q}_{I,J,K,n}+\frac{r}{n}\mathbf{t})\geq 0$ for all $(I,J,K)\in T_{r}^{n}$, $11\leq r\leq n-$ 1.
Proof.
For $j=1,\ldots,n$, define $\alpha_{n+1-j}$, $\beta_{n+1-j}$, $\gamma_{n+1-j}$ to be the respective constant values of the functions $Q_{1}$, $Q_{2}$, $Q_{3}$ on the interval $[(j-1)/n,j/n)$. The coordinates of $\alpha,\beta,\gamma$ are nonincreasing since the quantile functions $Q_{i}$ are nondecreasing. Now define
(3.10) $\hat{\alpha}_{i}:=-\alpha_{n+1-i},\quad\hat{\beta}_{i}:=-\beta_{n+1-i},\quad\text{and}\quad\hat{\gamma_{i}}:=-\gamma_{n+1-i}.$
The coordinates of $\hat{\alpha},\hat{\beta},\hat{\gamma}$ are also nonincreasing, and they are the eigenvalues of $n$-by-$n$ Hermitian matrices $A+B=C$ if and only if $\alpha,\beta,\gamma$ are. (This follows from considering the matrix equation $(-A)+(-B)=-C$.)
Now note that for $(I,J,K)\in T_{r}^{n}$, by Lemma 3.6 we have
$\frac{r}{n}\mathcal{E}\Big{(}\mathbf{Q},\mathbf{Q}_{I,J,K,n}+\frac{r}{n}\mathbf{t}\Big{)}$ $=-\sum_{i\in I}\int_{(i-1)/n}^{i/n}Q_{1}(t)\,\mathrm{d}t-\sum_{j\in J}\int_{(j-1)/n}^{j/n}Q_{2}(t)\,\mathrm{d}t+\sum_{k\in K}\int_{(k-1)/n}^{k/n}Q_{3}(t)\,\mathrm{d}t$
$=-\sum_{i\in I}\alpha_{n+1-i}-\sum_{j\in J}\beta_{n+1-j}+\sum_{k\in K}\gamma_{n+1-k}$
$=\sum_{i\in I}\hat{\alpha}_{i}+\sum_{j\in J}\hat{\beta}_{j}-\sum_{k\in K}\hat{\gamma}_{k}.$
In particular, we have $\mathcal{E}(\mathbf{Q},\mathbf{Q}_{I,J,K,n}+\frac{r}{n}\mathbf{t})\geq 0$ for all $(I,J,K)\in T_{r}^{n}$ if and only if $\sum_{k\in K}\hat{\gamma}_{k}\leq\sum_{i\in I}\hat{\alpha}_{i}+\sum_{j\in J}\hat{\beta}_{j}$ for all $(I,J,K)\in T_{r}^{n}$, which by Theorem 3.1 occurs if and only if $\hat{\alpha},\hat{\beta},\hat{\gamma}$ are the eigenvalues of Hermitian matrices $A+B=C$. In turn, this occurs if and only if $\alpha,\beta,\gamma$ are the eigenvalues of Hermitian matrices $A+B=C$. Since $\mathbf{Q}$ are the quantile functions corresponding to the spectra $\alpha,\beta,\gamma$, this completes the proof. ∎
3.4. An $n$-independent formulation of the Horn inequalities
In Proposition 3.7 we have stated the Horn inequalities for each fixed $n$ in terms of the quantile functions of empirical spectral measures. We now upgrade this to a uniform statement giving a single collection of infinitely many inequalities that characterises the solutions for all $n$ simultaneously.
The probability measure $\pi_{I,n}$ is obtained by considering $I$ as a subset of $\{1,\ldots,n\}$. Note that if $I\subset\{1,\ldots,n\}$, and $n^{\prime}\geq n$, we may also think of $I$ as a subset of $\{1,\ldots,n^{\prime}\}$. The relationship between the probability measures $\pi_{I,n}$ and $\pi_{I,n^{\prime}}$ can be expressed in terms of their quantile functions; namely, by (3.6) we have
(3.11) $Q_{I,n^{\prime}}(t)=\frac{n}{n^{\prime}}Q_{I,n}(t).$
Since $\mathcal{E}(\mathbf{Q},\mathbf{\tilde{Q}})$ is linear in its first argument, it follows in particular that for any sets $(I,J,K)$,
$\mathcal{E}(\mathbf{Q}_{I,J,K,n^{\prime}},\mathbf{Q}_{F,G,H,r})=\frac{n}{n^{\prime}}\mathcal{E}(\mathbf{Q}_{I,J,K,n},\mathbf{Q}_{F,G,H,r}),$
which in turn implies
$(I,J,K)\in T_{r}^{n}\implies(I,J,K)\in T_{r}^{n^{\prime}}\qquad\forall\,n^{\prime}\geq n.$
An alternative construction is by dilation. Namely, let $m\geq 1$ be an integer. With every subset $I$ of $\{1,\ldots,n\}$ of cardinality $r$ we may associate a subset $mI$ of $\{1,\ldots,mn\}$ of cardinality $mr$ by setting
$mI=\bigcup_{i\in I}\{mi-m+1,mi-m+2,\ldots,mi\}.$
A brief calculation tells us that with $\pi_{I,n}(\mathrm{d}x)$ as in (3.5) we have
$\pi_{mI,mn}=\pi_{I,n},\qquad\text{ or equivalently,}\qquad Q_{mI,mn}=Q_{I,n}.$
It follows in particular from Lemma 3.3 that
$U_{r}^{n}\subset U_{mr}^{mn}.$
In fact, we have the following lemma.
Lemma 3.8.
Let $I,J,K$ be a triple of subsets of $\{1,\ldots,n\}$ of the same cardinality $r$. Then
$(I,J,K)\in T_{r}^{n}\iff(mI,mJ,mK)\in T_{mr}^{mn}\qquad\text{for every integer $m\geq 1$}.$
Proof.
That the latter implies the former follows from setting $m=1$. On the other hand, if $(I,J,K)\in T_{r}^{n}$, then by Theorem 3.2, there are Hermitian matrices $A+B=C$ with eigenvalues given by the partitions $\lambda(I),\lambda(J),\lambda(K)$, or equivalently, with empirical spectra given by $\pi_{I,n},\pi_{J,n},\pi_{K,n}$ as in (3.5). By considering $mn$-by-$mn$ block diagonal matrices with $m$ copies of $n$-by-$n$ matrices $A$, $B$, and $C$, it follows that there exist $mn$-by-$mn$ Hermitian matrices with these same empirical spectra. However, since $(\pi_{I,n},\pi_{J,n},\pi_{K,n})=(\pi_{mI,mn},\pi_{mJ,mn},\pi_{mK,mn})$, it follows from the other direction of Theorem 3.2 that $(mI,mJ,mK)$ lies in $T_{mr}^{mn}$, as required. ∎
We now use the dilation property to prove the following forward property, which says that if a triple of quantile functions is constant on intervals of length $1/n$ and satisfies the Horn inequalities corresponding to all $(I,J,K)\in T_{r}^{n}$, then it satisfies the Horn inequalities corresponding to all $(I,J,K)$ in $T_{r^{\prime}}^{n^{\prime}}$ for all integers $r^{\prime},n^{\prime}$ with $11\leq r^{\prime}\leq n^{\prime}-$ 1 (c.f. the statement of Proposition 3.7). This is the desired $n$-independent formulation of the Horn inequalities: it shows that, for any value of $n$, a necessary and sufficient condition to be the spectra of a Hermitian triple is given by the full set of quantile function inequalities corresponding to all triples $(I,J,K)\in T^{n^{\prime}}_{r^{\prime}}$ for all values of $n^{\prime}$ and $r^{\prime}$.
Proposition 3.9.
Let $\mathbf{Q}$ be a triple of quantile functions that are constant on each interval $\left[\frac{j-1}{n},\frac{j}{n}\right)$ with $1\leq j\leq n$. The following two conditions are equivalent:
  1. (1)
    $\mathcal{E}\left(\mathbf{Q},\mathbf{Q}_{I,J,K,n}+\frac{r}{n}\mathbf{t}\right)\geq 0$ for all $(I,J,K)$ in all $T_{r}^{n}$, $11\leq r\leq n-$ 1.
  2. (2)
    $\mathcal{E}\left(\mathbf{Q},\mathbf{Q}_{I,J,K,n^{\prime}}+\frac{r^{\prime}}{n^{\prime}}\mathbf{t}\right)\geq 0$ for all $(I,J,K)$ in all $T_{r^{\prime}}^{n^{\prime}}$, where $r^{\prime},n^{\prime}$ are any integers $11\leq r^{\prime}\leq n^{\prime}-$ 1.
We remark that by Proposition 3.7, the property (1) is equivalent to $\mathbf{Q}$ being the spectral quantile functions of an $n$-by-$n$ Hermitian triple.
Proof.
It is clear that (2) implies (1).
To prove that (1) implies (2), fix some $(I,J,K)\in T_{r^{\prime}}^{n^{\prime}}$. We note that by Proposition 3.7, there are $n$-by-$n$ Hermitian matrices $A+B=C$ with spectral quantiles given by $\mathbf{Q}$. By considering block diagonal matrices (with $n^{\prime}$ copies of the same $n$-by-$n$ matrix along the diagonal), it is then easily verified that there are $n^{\prime}n$-by-$n^{\prime}n$ Hermitian matrices $A^{\prime}+B^{\prime}=C^{\prime}$ with these same empirical quantiles $\mathbf{Q}$. In particular, now using the other direction of Proposition 3.7, it follows that $\mathcal{E}\left(\mathbf{Q},\mathbf{Q}_{I,J,K,n^{\prime}n}+\frac{s}{n^{\prime}n}\mathbf{t}\right)\geq 0$ for any $(I,J,K)$ in $T_{s}^{n^{\prime}n}$ with $11\leq s\leq n^{\prime}n-$ 1.
Now by Lemma 3.8, $(I,J,K)\in T_{r^{\prime}}^{n^{\prime}}$ implies $(nI,nJ,nK)\in T_{r^{\prime}n}^{n^{\prime}n}$. Since there is a $n^{\prime}n$-by-$n^{\prime}n$ Hermitian triple with spectral quantiles $\mathbf{Q}$, it follows that
(3.12) $\mathcal{E}\left(\mathbf{Q},\mathbf{Q}_{nI,nJ,nK,nn^{\prime}}+\frac{r^{\prime}n}{n^{\prime}n}\mathbf{t}\right)\geq 0.$
However, since $\pi_{nI,nn^{\prime}}=\pi_{I,n^{\prime}}$, and likewise for $nJ$ and $nK$, it follows that
$\mathbf{Q}_{nI,nJ,nK,nn^{\prime}}+\frac{r^{\prime}n}{n^{\prime}n}\mathbf{t}=\mathbf{Q}_{I,J,K,n^{\prime}}+\frac{r^{\prime}}{n^{\prime}}\mathbf{t}.$
In particular, we have $\mathcal{E}\left(\mathbf{Q},\mathbf{Q}_{I,J,K,n^{\prime}}+\frac{r^{\prime}}{n^{\prime}}\mathbf{t}\right)\geq 0$, as required. ∎
3.5. Lattice and semilattice points
To close this section, we say a bit more about the geometry of the correspondence between spectra of Hermitian triples or subsets $(I,J,K)\in T^{n}_{r}$ and triples of quantile functions.
Definition 3.10 (Semilattice and lattice quantile functions).
  • Given an integer $r\geq 1$, we say that a quantile function $Q:[0,1]\to\mathbb{R}$ is $r$-semilattice if it is constant on each interval of the form $\left[\frac{j-1}{r},\frac{j}{r}\right)$, where $1\leq j\leq r$ is an integer.
  • Given another integer $n\geq 1$, we say that a quantile function $Q:[0,1]\to\mathbb{R}$ is $(n,r)$-lattice if it is $r$-semilattice and additionally, $Q(t)\in\frac{1}{n}\mathbb{Z}$ for all $t\in[0,1]$.
We say that a triple of quantile functions $\mathbf{Q}=(Q_{1},Q_{2},Q_{3})$ is $r$-semilattice or $(n,r)$-lattice if each $Q_{i}$ has the corresponding property.
Lemma 3.5, Proposition 3.7 and Proposition 3.9 together imply the following:
Corollary 3.11.
Characterisation of semilattice and lattice elements of $\mathscr{H}$ and $\mathscr{H}^{\prime}$:
  • An element $\mathbf{Q}\in\mathscr{H}^{\prime}$ is $r$-semilattice if and only if $\mathbf{Q}$ are the empirical quantiles of an $r$-by-$r$ Hermitian triple.
  • An element $\mathbf{Q}\in\mathscr{H}$ satisfies $\mathbf{Q}=\mathbf{Q}_{I,J,K,n}$ for some $(I,J,K)\in T_{r}^{n}$ if and only if $\mathbf{Q}$ is $(n,r)$-lattice and each $Q_{i}$ takes values in $[0,1-r/n]$.
From the finite-$n$ Horn inequalities it is then clear that, to determine whether a triple of $n$-semilattice quantile functions belongs to $\mathscr{H}^{\prime}$, one need only consider inequalities corresponding to $(n,r)$-lattice elements of $\mathscr{H}$ taking values in $[0,1-r/n]$. Together with Theorem 2.3, this implies:
Corollary 3.12.
Let $\mathbf{Q}$ be a triple of $n$-semilattice quantile functions. Then the following are equivalent:
  1. (1)
    $\mathcal{E}(\mathbf{Q},\mathbf{\tilde{Q}}+\mu\mathbf{t})\geq 0$ for every $\mathbf{\tilde{Q}}\in\mathscr{H}$ and $0\leq\mu\leq\eta_{\mathbf{\tilde{Q}}}$.
  2. (2)
    $\mathcal{E}(\mathbf{Q},\mathbf{\tilde{Q}}+\frac{r}{n}\mathbf{t})\geq 0$ for every $(n,r)$-lattice element $\mathbf{\tilde{Q}}$ of $\mathscr{H}$ taking values in $[0,1-r/n]$, for all $11\leq r\leq n-$ 1.
4. Proofs of the main results
We now turn to proving our main results. First we study the continuity properties of the composition functional $\mathcal{E}$. Although $\mathcal{E}$ is not actually continuous in either argument in our chosen topology, it turns out to have enough continuity for our purposes: it is continuous in its first argument assuming that the second argument is a triple of quantile functions of measures with bounded densities, while in its second argument it is continuous along sequences of triples with uniformly bounded densities. We then prove Theorem 4.4, which gives a quantitative bound on the distance between points of $\mathscr{H}$ and triples representing Horn inequalities in specific sets $T^{n}_{r}$, and we use it to prove both Theorem 2.1 (on approximation by Horn inequalities with fixed asymptotic ratio $r_{n}/n\to q$) and Theorem 4.7 (on redundancy in the infinite system of Horn inequalities). Finally, we prove Theorem 2.3 on self-characterisation.
4.1. Continuity properties of the composition functional
Recall that a sequence of probability measures $(\pi^{n})_{n\geq 1}$ on $\mathbb{R}$ are said to converge weakly to $\pi$ if for all bounded and continuous functions $f:\mathbb{R}\to\mathbb{R}$ we have
$\int_{-\infty}^{\infty}f(x)\,\pi^{n}(\mathrm{d}x)\to\int_{-\infty}^{\infty}f(x)\,\pi(\mathrm{d}x).$
In this case we write $\pi^{n}\Longrightarrow\pi$. Note that if $(Q^{n})_{n\geq 1}$ and $Q$ are the associated quantile functions, then by (2.4), weak convergence is equivalent to
$\int_{0}^{1}f(Q^{n}(t))\,\mathrm{d}t\to\int_{0}^{1}f(Q(t))\,\mathrm{d}t.$
Another equivalent definition is that $\pi^{n}\Longrightarrow\pi$ if and only if $Q^{n}(t)\to Q(t)$ at every continuity point of $Q(t)$ [17, Prop. 5, p. 250].
Recall further that we have topologised the space of finite-mean probability measures using the Wasserstein distance, and that Wasserstein convergence is equivalent to weak convergence plus convergence in expectation, or to $L^{1}$ convergence of quantile functions. For measures with uniformly bounded support, weak convergence implies convergence in Wasserstein distance.
Given triples $\mathbf{Q}^{n}=(Q^{n}_{1}(t),Q^{n}_{2}(t),Q^{n}_{3}(t))$ and $\mathbf{Q}=(Q_{1}(t),Q_{2}(t),Q_{3}(t))$ of quantile functions, we say that $\mathbf{Q}^{n}$ converges weakly (resp. in Wasserstein distance) to $\mathbf{Q}$ if each $Q_{i}^{n}$ converges to $Q_{i}$ pointwise at points of continuity of $Q_{i}$ (resp. in $L^{1}$).
Suppose that a probability measure $\pi$ has a density with respect to Lebesgue measure that is bounded above by $L>0$. It is not hard to see that this is equivalent to the property that its quantile function satisfies $Q(s+t)-Q(t)\geq s/L$ for all $0\leq t\leq s+t\leq 1$.
Lemma 4.1.
Let $\mathbf{P}$ be a triple of quantile functions in $L^{1}([0,1])$. Let $(\mathbf{Q}^{n})_{n\geq 1}$ be a sequence of triples of quantile functions associated with triples $(\bm{\pi}^{n})_{n\geq 1}$ of probability measures supported on $[0,1]$ with density functions uniformly bounded above by some $K>0$, and suppose that $\bm{\pi}^{n}$ converges weakly to $\bm{\pi}$. Then
(4.1) $\lim_{n\to\infty}\mathcal{E}(\mathbf{P},\mathbf{Q}^{n})=\mathcal{E}(\mathbf{P},\mathbf{Q}).$
Alternatively, let $\mathbf{P}$ be a triple of quantile functions in $L^{1}([0,1])$, and let $\mathbf{Q}$ be a triple of quantile functions of probability measures supported on $[0,1]$ with bounded densities. Let $\mathbf{P}^{n}$ be a sequence of triples converging to $\mathbf{P}$ in Wasserstein distance. Then
(4.2) $\lim_{n\to\infty}\mathcal{E}(\mathbf{P}^{n},\mathbf{Q})=\mathcal{E}(\mathbf{P},\mathbf{Q}).$
Proof.
Write $\mathcal{E}(\mathbf{P},\mathbf{Q}):=-E(P_{1},Q_{1})-E(P_{2},Q_{2})+E(P_{3},Q_{3})$ where
(4.3) $E(P,Q):=\int_{0}^{1}P(Q(t))\mathrm{d}t.$
It is then sufficient to study the continuity properties of $E(P,Q)$.
We now prove (4.1). It is sufficient to show that whenever $Q_{n}(t)$ is a sequence of quantile functions satisfying $Q_{n}(t+s)-Q_{n}(t)\geq s/K$ for all $0\leq t\leq s+t\leq 1$, and such that $Q_{n}(t)\to Q(t)$ at every continuity point $t$ of $Q(t)$, we have $E(P,Q_{n})\to E(P,Q)$.
Since the continuous functions are dense in $L^{1}([0,1])$, for any $\varepsilon>0$ we may choose a continuous $f:[0,1]\to[0,1]$ such that $\int_{0}^{1}|f(t)-P(t)|\mathrm{d}t\leq\varepsilon$. Now choose $n_{0}$ sufficiently large such that for all $n\geq n_{0}$ we have $\left|\int_{0}^{1}f(Q_{n}(t))\mathrm{d}t-\int_{0}^{1}f(Q(t))\mathrm{d}t\right|\leq\varepsilon$. Then by the triangle inequality,
$\left|\int_{0}^{1}P(Q_{n}(t))-P(Q(t))\mathrm{d}t\right|$ $\leq\left|\int_{0}^{1}P(Q_{n}(t))-f(Q_{n}(t))\mathrm{d}t\right|+\left|\int_{0}^{1}f(Q_{n}(t))-f(Q(t))\mathrm{d}t\right|$
(4.4) $+\left|\int_{0}^{1}P(Q(t))-f(Q(t))\mathrm{d}t\right|.$
Now by construction, for all $n\geq n_{0}$ we have $\left|\int_{0}^{1}f(Q_{n}(t))\mathrm{d}t-\int_{0}^{1}f(Q(t))\mathrm{d}t\right|\leq\varepsilon$. Moreover, if $Q(t)$ is the quantile function of a measure $\pi$ with density bounded above by $K$, we have
(4.5) $\left|\int_{0}^{1}P(Q(t))-f(Q(t))\mathrm{d}t\right|\leq\int_{0}^{1}|P(x)-f(x)|\pi(\mathrm{d}x)\leq K\int_{0}^{1}|P(x)-f(x)|\mathrm{d}x\leq K\varepsilon.$
The same bound as (4.5) holds with $Q_{n}(t)$ in place of $Q(t)$, and thus using (4.1), for all $n\geq n_{0}$ we have
$\left|\int_{0}^{1}P(Q_{n}(t))-P(Q(t))\mathrm{d}t\right|\leq(2K+1)\varepsilon.$
Since $\varepsilon$ is arbitrary, the proof of (4.1) follows.
We turn to proving (4.2). Let $Q(t)$ be the quantile function of a probability measure $\pi$ that is supported on $[0,1]$ and has a bounded density with respect to Lebesgue measure. Then $E(P,Q):=\int_{0}^{1}P(Q(t))\mathrm{d}t=\int_{0}^{1}P(x)\pi(\mathrm{d}x)=\int_{0}^{1}P(x)\rho(x)\mathrm{d}x$ for some bounded function $\rho$. If $P^{n}\to P$ in $L^{1}$ then $P^{n}\rho\to P\rho$ in $L^{1}$, thus $E(P^{n},Q)\to E(P,Q)$, completing the proof of (4.2). ∎
In particular, we have the following corollary.
Corollary 4.2.
The trace functional is continuous in the Wasserstein metric. That is, if $\mathbf{P}^{n}\to\mathbf{P}$ in Wasserstein distance, then
$\lim_{n\to\infty}\mathrm{tr}(\mathbf{P}^{n})\to\mathrm{tr}(\mathbf{P}).$
Proof.
Recall that $\mathcal{E}(\mathbf{P},\mathbf{t})=\mathrm{tr}(\mathbf{P})$. Set $\mathbf{Q}=\mathbf{t}$ in (4.2). ∎
We will require one further lemma on specific limits of $\mathcal{E}$.
Lemma 4.3.
Let $\mathbf{Q}$ be a triple of bounded quantile functions and let $\mathbf{\tilde{Q}}$ be a triple of quantile functions taking values in the half-open interval $[0,1)$. Then
(4.6) $\lim_{\varepsilon\to 0^{+}}\mathcal{E}(\mathbf{Q},\mathbf{\tilde{Q}}+\varepsilon\mathbf{t})=\mathcal{E}(\mathbf{Q},\mathbf{\tilde{Q}}).$
Proof.
The assumption that the quantile functions in $\mathbf{\tilde{Q}}$ take values in $[0,1)$ is required to ensure that $\mathcal{E}(\mathbf{Q},\mathbf{\tilde{Q}}+\varepsilon\mathbf{t})$ is well defined for sufficiently small $\varepsilon>0$. Then it is enough to show that $E(Q_{i},\tilde{Q}_{i}+\varepsilon t)\to E(Q_{i},\tilde{Q}_{i})$ for $E$ defined as in (4.3). Since $Q_{i}$ is bounded and right-continuous with compact domain, it is uniformly right-continuous; thus for every $\delta>0$ and for $\varepsilon>0$ sufficiently small, we have $|Q_{i}(\tilde{Q}_{i}(t)+\varepsilon t)-Q_{i}(\tilde{Q}_{i}(t))|<\delta$, and thus
$|E(Q_{i},\tilde{Q}_{i}+\varepsilon t)-E(Q_{i},\tilde{Q}_{i})|\leq\int_{0}^{1}|Q_{i}(\tilde{Q}_{i}(t)+\varepsilon t)-Q_{i}(\tilde{Q}_{i}(t))|\,\mathrm{d}t<\delta,$
which proves the claim. ∎
4.2. Approximation by Horn inequalities
Recall that $\mathscr{H}$ is the topological closure of the set of triples of quantile functions of $n$-by-$n$ Hermitian matrices with eigenvalues in $[0,1]$. The following theorem says not only that triples of the form $\mathbf{Q}_{I,J,K,n}$ associated with $(I,J,K)$ in $T_{r}^{n}$ are dense in $\mathscr{H}$ as $n$ and $r$ range over all values, but also that any given element $\mathbf{Q}$ of $\mathscr{H}$ may be approximated by a sequence $(I_{n},J_{n},K_{n})\in T_{r_{n}}^{n}$ with $r_{n}/n$ converging to a prescribed ratio. This is our first main result.
See 2.1
In fact, we will deduce Theorem 2.1 from the following more quantitative statement.
Theorem 4.4.
Let $\mathbf{Q}\in\mathscr{H}[0,1-\mu)$ for $\mu\in(0,1)$, and let $\delta\in(0,\mu)$. Then for any pair $(n,r)$ with $136/\delta\leq r\leq n-$ 1, there exists an $(n,r)$-lattice triple $\mathbf{Q}^{\prime}\in\mathscr{H}[0,1-\mu+\delta)$ that satisfies
$\int_{0}^{1}|Q_{i}^{\prime}(t)-Q_{i}(t)|\,\mathrm{d}t<\delta.$
If it is also the case that $r\leq(\mu-\delta)n$, then $\mathbf{Q}^{\prime}=\mathbf{Q}_{I,J,K,n}$ for some $(I,J,K)\in T_{r}^{n}$.
Before giving the proof of Theorem 4.4, we state and prove two lemmas.
Lemma 4.5.
Let $\mathbf{Q}$ be an $r$-semilattice triple in $\mathscr{H}[0,1-\mu)$ for $\mu>0$. Then for every $\delta\in(0,\mu)$ and every $n\geq 18/\delta$, there exists an $(n,r)$-lattice triple $\mathbf{Q}^{\prime}\in\mathscr{H}[0,1-\mu+\delta)$ approximating $\mathbf{Q}$ in the sense that
$|Q^{\prime}_{i}(t)-Q_{i}(t)|\leq\delta\qquad\text{for all $t\in[0,1]$.}$
Proof.
We would like to approximate $\mathbf{Q}$ by an $(n,r)$-lattice triple in $\mathscr{H}[0,1-\mu+\delta)$, but first we obtain some leeway by adding a small perturbation to $\mathbf{Q}$ so as to ensure that the relevant Horn inequalities hold in a way uniformly bounded away from zero. To this end, we define a new triple $\mathbf{S}=(S_{1}(t),S_{2}(t),S_{3}(t))$ by setting
(4.7) $S_{i}(t):=s+\mathbb{1}_{i=3}\frac{r+1}{2}\qquad t\in\left[\frac{s-1}{r},\frac{s}{r}\right).$
It is easily verified that $\mathrm{tr}(\mathbf{S})=0$, and we will show in a moment that $\mathbf{S}$ is an element of $\mathscr{H}$. In the meantime define a new triple $\mathbf{Q}^{(\varepsilon)}$ by setting
$\mathbf{Q}^{(\varepsilon)}:=\mathbf{Q}+\varepsilon\mathbf{S}.$
Noting that $20\leq S_{i}(t)\leq(3r+1)/$ 2 for all $t\in[0,1]$ and $i=1,2,3$, we have
(4.8) $|Q_{i}^{(\varepsilon)}(t)-Q_{i}(t)|\leq(3r+1)\varepsilon/2\leq 2r\varepsilon\qquad\text{for all $t\in[0,1],\ i=1,2,3$}.$
In particular, $Q_{i}^{(\varepsilon)}$ takes values in $[0,1-\mu+2r\varepsilon)$.
For sufficiently small $\varepsilon$, $\mathbf{Q}^{(\varepsilon)}$ takes values in $[0,1-\mu)$ as well. Also, for $(F,G,H)\in T_{p}^{r}$, by linearity we have
$\mathcal{E}\Big{(}\mathbf{Q}^{(\varepsilon)},\mathbf{Q}_{F,G,H,r}+\frac{p}{r}\mathbf{t}\Big{)}$ $=\mathcal{E}\Big{(}\mathbf{Q},\mathbf{Q}_{F,G,H,r}+\frac{p}{r}\mathbf{t}\Big{)}+\varepsilon\mathcal{E}\left(\mathbf{S},\mathbf{Q}_{F,G,H,r}+\frac{p}{r}\mathbf{t}\right)$
$\geq\varepsilon\mathcal{E}\left(\mathbf{S},\mathbf{Q}_{F,G,H,r}+\frac{p}{r}\mathbf{t}\right),$
where to obtain the final inequality above we have used the fact that, since $\mathbf{Q}$ are the spectral quantiles of an $r$-by-$r$ Hermitian triple, by Proposition 3.7 we have $\mathcal{E}(\mathbf{Q},\mathbf{Q}_{F,G,H,r}+\frac{p}{r}\mathbf{t})\geq 0$ for all $(F,G,H)\in T_{p}^{r}$.
Now we lower bound $\mathcal{E}\left(\mathbf{S},\mathbf{Q}_{F,G,H,r}+\frac{p}{r}\mathbf{t}\right)$. Using Lemma 3.6 to obtain the second equality below, and then (4.7) to obtain the third, we have
$\mathcal{E}\left(\mathbf{S},\mathbf{Q}_{F,G,H,r}+\frac{p}{r}\mathbf{t}\right)$ $=-\int_{0}^{1}S_{1}(Q_{F,r}(t)+\frac{p}{r}t)\mathrm{d}t-\int_{0}^{1}S_{2}(Q_{G,r}(t)+\frac{p}{r}t)\mathrm{d}t$
$\qquad+\int_{0}^{1}S_{3}(Q_{H,r}(t)+\frac{p}{r}t)\mathrm{d}t$
$=\frac{r}{p}\left(-\sum_{f\in F}\int_{(f-1)/r}^{f/r}S_{1}(t)\mathrm{d}t-\sum_{g\in G}\int_{(g-1)/r}^{g/r}S_{2}(t)\mathrm{d}t+\sum_{h\in H}\int_{(h-1)/r}^{h/r}S_{3}(t)\mathrm{d}t\right)$
$=\frac{r}{p}\left(-\sum_{f\in F}f-\sum_{g\in G}g+\sum_{h\in H}\left(h+\frac{r+1}{2}\right)\right).$
For $(F,G,H)\in T_{p}^{r}\subset U_{p}^{r}$, the definition (3.2) implies that
$-\sum_{f\in F}f-\sum_{g\in G}g+\sum_{h\in H}h=-\frac{p(p+1)}{2},$
and so, together with the fact that $H$ has cardinality $p$, we have
$\mathcal{E}\left(\mathbf{S},\mathbf{Q}_{F,G,H,r}+\frac{p}{r}\mathbf{t}\right)=\frac{r}{p}\left(-\frac{p(p+1)}{2}+p\frac{r+1}{2}\right)=\frac{r(r-p)}{2}.$
In particular, since $11\leq p\leq r-$ 1, it follows that for every $p$ and for every $(F,G,H)\in T_{p}^{r}$ we have
$\mathcal{E}\Big{(}\mathbf{Q}^{(\varepsilon)},\mathbf{Q}_{F,G,H,r}+\frac{p}{r}\mathbf{t}\Big{)}\geq\varepsilon r/2>0.$
Thus $\mathbf{Q}^{(\varepsilon)}=(Q_{1}^{(\varepsilon)}(t),Q_{2}^{(\varepsilon)}(t),Q_{3}^{(\varepsilon)}(t))$ satisfies the relevant Horn inequalities in a way uniformly bounded away from zero.
For all sufficiently large $n$, we now construct a triple $\mathbf{Q}^{\prime}=\mathbf{Q}^{(\varepsilon,n)}\in\mathscr{H}[0,1-\mu+\delta)$ that satisfies the conditions of the theorem. In this direction, given any integer $n\geq 1$, we begin by defining an auxillary triple $\mathbf{P}^{(\varepsilon,n)}$ by setting
$P_{i}^{(\varepsilon,n)}(t):=\text{Largest multiple of $1/n$ lessthan $Q_{i}^{(\varepsilon)}(t)$}.$
Then $\mathbf{P}^{(\varepsilon,n)}:=(P_{i}^{(\varepsilon,n)}(t),P_{2}^{(\varepsilon,n)}(t),P_{3}^{(\varepsilon,n)}(t))$ is a triple of quantile functions satisfying
(4.9) $P_{i}^{(\varepsilon,n)}(t)\leq Q_{i}^{(\varepsilon)}(t)\leq P_{i}^{(\varepsilon,n)}(t)+1/n$
for each $t\in[0,1]$ and $i=1,2,3$. However, in general $\mathrm{tr}(\mathbf{P}^{(\varepsilon,n)})\neq 0$. The remainder of the proof is dedicated to remedying this issue.
By (4.9) and the fact that $\mathrm{tr}(\mathbf{Q}^{(\varepsilon)})=0$ we have
(4.10) $-1/n\leq\mathrm{tr}(\mathbf{P}^{(\varepsilon,n)})\leq 2/n.$
Also, since each $P_{i}^{(\varepsilon,n)}(t)$ is constant on intervals of length $1/r$ and takes values in integer multiples of $1/n$, it follows that $\mathrm{tr}(\mathbf{P}^{(\varepsilon,n)})=\tau_{\varepsilon,n}/nr$ for some integer $\tau_{\varepsilon,n}$. By (4.10), we have $-r\leq\tau_{\varepsilon,n}\leq 2r$. Now choose three integers $0\leq a_{1},a_{2},a_{3}\leq r$ such that $a_{1}+a_{2}-a_{3}=\tau_{\varepsilon,n}$, and define
$\mathbf{Q}^{(\varepsilon,n)}_{i}(t):=\mathbf{P}^{(\varepsilon,n)}_{i}(t)+\frac{1}{n}\mathbb{1}\left\{t\geq\frac{r-a_{i}}{r}\right\}.$
Then $\mathrm{tr}(\mathbf{Q}^{(\varepsilon,n)})=\mathrm{tr}(\mathbf{P}^{(\varepsilon,n)})+\frac{1}{nr}(a_{1}+a_{2}-a_{3})=0$ by the definitions of $a_{1},a_{2},a_{3}$. Note also that $|Q_{i}^{(\varepsilon,n)}(t)-Q_{i}^{(\varepsilon)}(t)|\leq 1/n$, and in combination with (4.8) this implies that
(4.11) $|Q_{i}^{(\varepsilon,n)}(t)-Q_{i}(t)|\leq 2r\varepsilon+1/n\qquad\text{for all $t\in[0,1],\ i=1,2,3$}.$
In particular, $Q_{i}^{(\varepsilon,n)}$ takes values in $[0,1-\mu+2r\varepsilon+1/n)$.
We have thus far constructed a trace-zero triple $\mathbf{Q}^{(\varepsilon,n)}$ taking values in multiples of $1/n$ and approximating $\mathbf{Q}$ in the sense that (4.11) holds. We now verify that $\mathbf{Q}^{(\varepsilon,n)}$ is also an element of $\mathscr{H}$ for $n$ sufficiently large.
To this end, by Proposition 3.7, we need to establish that
$\mathcal{E}\Big{(}\mathbf{Q}^{(\varepsilon,n)},\mathbf{Q}_{F,G,H,r}+\frac{p}{r}\mathbf{t}\Big{)}\geq 0$
for all $11\leq p\leq r-$ 1 and all $(F,G,H)\in T_{p}^{r}$. This indeed holds provided $n\geq 6/\varepsilon r$: since $|Q^{(\varepsilon,n)}_{i}(t)-Q^{(\varepsilon)}_{i}(t)|\leq 1/n$ for all $i=1,2,3$ and all $t\in[0,1]$, we then have
$\mathcal{E}\Big{(}\mathbf{Q}^{(\varepsilon,n)},\mathbf{Q}_{F,G,H,r}+\frac{p}{r}\mathbf{t}\Big{)}\geq\mathcal{E}\Big{(}\mathbf{Q}^{(\varepsilon)},\mathbf{Q}_{F,G,H,r}+\frac{p}{r}\mathbf{t}\Big{)}-3/n\geq\varepsilon r/2-3/n\geq 0,$
where to obtain the final inequality above, we used $n\geq 6/\varepsilon r$.
We now take stock of what we have achieved so far. We began with an $r$-semilattice triple $\mathbf{Q}\in\mathscr{H}[0,1-\mu)$. Note that $2r\varepsilon+1/n\leq 3r\varepsilon$ whenever $n\geq 6/r\varepsilon$, and it follows that for any $\varepsilon>0$ and any $n\geq 6/\varepsilon r$ we have constructed an $(n,r)$-lattice triple in $\mathscr{H}[0,1-\mu+3r\varepsilon)$ satisfying
(4.12) $|Q_{i}^{(\varepsilon,n)}(t)-Q_{i}(t)|\leq 3r\varepsilon\qquad\text{for all $t\in[0,1],\ i=1,2,3$}.$
Since $\varepsilon>0$ was arbitrary, we can obtain the result as written by choosing $\varepsilon$ so that $\delta=3r\varepsilon$. ∎
Lemma 4.6.
Given a quantile function $Q:[0,1]\to[0,1]$, define its $r$-semilattice average by
$Q^{r}(t)=r\int_{(j-1)/r}^{j/r}Q(u)\,\mathrm{d}u\qquad\text{for $t\in\left[\frac{j-1}{r},\frac{j}{r}\right),\quad 1\leq j\leq r$.}$
Then
$||Q^{r}-Q||_{1}=\int_{0}^{1}|Q^{r}(t)-Q(t)|\,\mathrm{d}t\leq\frac{1}{r}.$
Proof.
Let $a_{j}:=Q(j/r)-Q((j-1)/r)$. Then $\sum_{j=1}^{r}a_{j}\leq 1$. Moreover,
$\left|\int_{(j-1)/r}^{j/r}Q^{r}(u)-Q(u)\,\mathrm{d}u\right|\leq\frac{a_{j}}{r}.$
The result now follows. ∎
With the preceding groundwork, Theorems 4.4 and 2.1 are quick to prove.
Proof of Theorem 4.4.
Let $\mathbf{Q}$ be any triple of quantile functions in $\mathscr{H}[0,1-\mu)$. Let $n,r$ be any integers with $136/\delta\leq r\leq n-$ 1. Let $\mathbf{Q}^{r}$ be the $r$-semilattice average of $\mathbf{Q}$, so that $||Q^{r}_{i}-Q_{i}||_{1}\leq 1/r$ by Lemma 4.6. Note that the quantile functions in the triple $\mathbf{Q}^{r}$ also take values in $[0,1-\mu)$. Let $\sigma=\delta/2$. Note that since $n>r\geq 36/\delta=18/\sigma$, by Lemma 4.5 there exists an $(n,r)$-lattice triple $\mathbf{Q}^{\prime}\in\mathscr{H}[0,1-\mu+\sigma)$ approximating $\mathbf{Q}^{r}$ in that $|Q^{\prime}_{i}(t)-Q_{i}^{r}(t)|\leq\sigma$ for all $t\in[0,1]$ and $i=1,2,3$, which in particular implies that $||Q_{i}^{\prime}-Q_{i}^{r}||_{1}\leq\sigma$. It now follows from the triangle inequality that
$||Q_{i}-Q_{i}^{\prime}||_{1}\leq||Q_{i}-Q_{i}^{r}||+||Q_{i}^{r}-Q_{i}^{\prime}||\leq 1/r+\sigma\leq\delta/36+\delta/2<\delta,$
which proves the desired inequality.
If $r\leq(\mu-\delta)n$, then $1-\mu+\delta\leq 1-r/n$, so that each $Q_{i}^{\prime}$ is $(n,r)$-lattice and takes values in $[1,1-r/n)$. Then by Corollary 3.11, $\mathbf{Q}^{\prime}=\mathbf{Q}_{I,J,K,n}$ for some $(I,J,K)\in T_{r}^{n}$. ∎
Proof of Theorem 2.1.
It suffices to prove the statement assuming that $\mathbf{Q}\in\mathscr{H}[0,1-\mu)$, $\mu\in(0,1)$, and $0<q<\mu$. The full statement then follows by a diagonal argument, using the dilation invariance 2.3.2 to approximate triples taking values in $[0,1-\mu]$ with triples taking values in $[0,1-\mu)$, then approximating again to extend to $0\leq q\leq\mu$ and finally $0\leq\mu\leq 1$.
Under these assumptions, choose a sequence $r_{n}$ such that $r_{n}/n\to q\in(0,\mu)$. For $\delta$ sufficiently small and $n$ sufficiently large depending on $\delta$, we have $r_{n}\geq 36/\delta$ and $r_{n}\leq(\mu-\delta)n$. Then by Theorem 4.4, there exists $\mathbf{Q}^{\prime}=\mathbf{Q}_{I,J,K,n}$ for some $(I,J,K)\in T_{r_{n}}^{n}$ such that $\|Q_{i}^{\prime}-Q_{i}\|_{1}\leq\delta$, and this completes the proof. ∎
These approximation results allow us to identify many sparse subsets of Horn inequalities that are equivalent to the entire system in the sense that they have the same set of solutions. The following theorem is our second main result and was stated informally above as Theorem 2.2.
Theorem 4.7.
Choose any sequences $(n_{k})_{k\geq 1}$, $(r_{k})_{k\geq 1}$ of positive integers such that $(r_{k}/n_{k})_{k\geq 1}$ is a dense subset of $(0,1)$. Let $\mathbf{Q}$ be a triple of quantile functions in $L^{1}([0,1])$ satisfying $\mathrm{tr}(\mathbf{Q})=0$. Then $\mathbf{Q}\in\mathscr{H}^{\prime}$ if and only if
(4.13) $\mathcal{E}\Big{(}\mathbf{Q},\mathbf{Q}_{I,J,K,n_{k}}+\frac{r_{k}}{n_{k}}\mathbf{t}\Big{)}\geq 0\quad\text{ for all $(I,J,K)\in T^{n_{k}}_{r_{k}}$, $\,k\geq 1$.}$
In other words, the finite-dimensional Horn inequalities indexed by triples in the sets $(T^{n_{k}}_{r_{k}})_{k\geq 1}$ imply the full set of Horn inequalities.
Proof.
The “only if” direction is immediate from Theorem 2.3, which we prove below.
For the “if” direction, we show that if $\mathbf{Q}^{n}$ is the $n$-average of the triple $\mathbf{Q}$, then (4.13) entails that $\mathbf{Q}^{n}$ lies in $\mathcal{H}^{\prime}$ for every $n$. Since $\mathcal{H}^{\prime}$ is closed and $\mathbf{Q}^{n}$ converges to $\mathbf{Q}$, this will tell us that $\mathbf{Q}$ lies in $\mathcal{H}^{\prime}$ also. In this direction, we begin by showing that if the inequalities (4.13) hold, then for $n$ sufficiently large,
$\mathcal{E}\Big{(}\mathbf{Q}^{n},\mathbf{Q}_{I,J,K,n}+\frac{r}{n}\mathbf{t}\Big{)}\geq 0\quad\text{ for all $(I,J,K)\in T^{n}_{r}$, $1\leq r\leq n-1$.}$
Note that the assumptions imply that $n_{k}>r_{k}$ and that $r_{k}$ is unbounded. By Theorem 4.4, for any $(I,J,K)\in T^{n}_{r}$ we can choose subsequences $(n_{k_{j}})_{j\geq 1}$, $(r_{k_{j}})_{j\geq 1}$ and a sequence $(F_{j},G_{j},K_{j})\in T^{n_{k_{j}}}_{r_{k_{j}}}$ such that $r_{k_{j}}/n_{k_{j}}\to r/n$ and
$\mathbf{Q}_{F_{j},G_{j},K_{j},n_{k_{j}}}+\frac{r_{k_{j}}}{n_{k_{j}}}\mathbf{t}\ \ \longrightarrow\ \ \mathbf{Q}_{I,J,K,n}+\frac{r}{n}\mathbf{t}.$
Since the quantile functions in the sequence of triples above correspond to measures with densities that are eventually bounded above by $2n/r$, by Lemma 4.1 we have
$0\leq\mathcal{E}\Big{(}\mathbf{Q}^{n},\mathbf{Q}_{F_{j},G_{j},K_{j},n_{k_{j}}}+\frac{r_{k_{j}}}{n_{k_{j}}}\mathbf{t}\Big{)}\ \ \longrightarrow\ \ \mathcal{E}\Big{(}\mathbf{Q}^{n},\mathbf{Q}_{I,J,K,n}+\frac{r}{n}\mathbf{t}\Big{)}\geq 0$
Therefore, by Proposition 3.7, $\mathbf{Q}^{n}$ are the spectral quantiles of Hermitian matrices $A+B=C$, while by Lemma 4.6, $\mathbf{Q}^{n}\to\mathbf{Q}$, and thus $\mathbf{Q}\in\mathscr{H}^{\prime}$. ∎
The inequalities (4.13) correspond to actual Horn inequalities for $n_{k}$-by-$n_{k}$ matrices, and the requirement that the values of $r_{k}/n_{k}$ be dense in $(0,1)$ merely ensures that we can approximate the shifted triples $\mathbf{Q}_{I,J,K,n}+\frac{r}{n}\mathbf{t}$ corresponding to all of the other Horn inequalities. On the other hand, if we allow ourselves to shift triples of quantile functions by arbitrary multiples of $\mathbf{t}$ as in Theorem 2.3, then we can recover the entire system only from triples for which $r_{k}/n_{k}\to 0$ arbitrarily quickly.
Corollary 4.8.
Choose any sequences $(n_{k})_{k\geq 1}$, $(r_{k})_{k\geq 1}$ of positive integers such that $r_{k}<n_{k}$, $r_{k}\to\infty$, and $r_{k}/n_{k}\to 0$. Let $\mathbf{Q}$ be a triple of quantile functions in $L^{1}([0,1])$ satisfying $\mathrm{tr}(\mathbf{Q})=0$. Then $\mathbf{Q}\in\mathscr{H}^{\prime}$ if and only if
$\mathcal{E}\big{(}\mathbf{Q},\mathbf{Q}_{I,J,K,n_{k}}+\mu\mathbf{t}\big{)}\geq 0\quad\text{ for all $(I,J,K)\in T^{n_{k}}_{r_{k}}$, $k\geq 1$, and $0<\mu\leq\eta_{\mathbf{Q}_{I,J,K,n_{k}}}$.}$
The proof is very similar to that of Theorem 4.7, and we omit it.
4.3. The self-averaging and self-characterisation properties
Next, we prove the self-averaging property stated in Section 2.2.
See 2.8
Proof.
We first point out that each function $Q_{i}^{n}$ is indeed a valid quantile function of a probability measure with bounded support. It is right-continuous by construction, it is nonincreasing because $Q_{i}$ is nonincreasing, and it is bounded because $Q_{i}\in L^{1}$. Moreover $\operatorname{tr}(\mathbf{Q}^{n})=\operatorname{tr}(\mathbf{Q})=0$.
We now claim that for each $n\geq 1$, $\mathbf{Q}^{n}$ are the spectral quantile functions of a triple of $n$-by-$n$ Hermitian matrices $A+B=C$. To see this, by Proposition 3.7 we need to verify that $\mathcal{E}(\mathbf{Q}^{n},\mathbf{Q}_{I,J,K,n}+\frac{r}{n}\mathbf{t})\geq 0$ for all $(I,J,K)\in T_{r}^{n}$, $11\leq r\leq n-$ 1. However, we actually claim that for all $(I,J,K)\in T_{r}^{n}$, we have the equality
(4.14) $\mathcal{E}\Big{(}\mathbf{Q}^{n},\mathbf{Q}_{I,J,K,n}+\frac{r}{n}\mathbf{t}\Big{)}=\mathcal{E}\Big{(}\mathbf{Q},\mathbf{Q}_{I,J,K,n}+\frac{r}{n}\mathbf{t}\Big{)}.$
To see that (4.14) holds, first note that the definition (2.12) of $Q_{i}^{n}$ entails
(4.15) $\int_{(j-1)/n}^{j/n}Q_{i}^{n}(t)\,\mathrm{d}t=\int_{(j-1)/n}^{j/n}Q_{i}(t)\,\mathrm{d}t$
for all $1\leq j\leq n$. Now using Lemma 3.6 to obtain the first and third equalities below, and (4.15) to obtain the second equality, for any subset $I\subset\{1,\ldots,n\}$ of cardinality $r$ we have
$\int_{0}^{1}Q_{i}^{n}\left(Q_{I,n}(t)+\frac{r}{n}t\right)\mathrm{d}t$ $=\frac{n}{r}\sum_{i\in I}\int_{(i-1)/n}^{i/n}Q_{i}^{n}(t)\,\mathrm{d}t$
$=\frac{n}{r}\sum_{i\in I}\int_{(i-1)/n}^{i/n}Q_{i}(t)\,\mathrm{d}t$
(4.16) $=\int_{0}^{1}Q_{i}\left(Q_{I,n}(t)+\frac{r}{n}t\right)\mathrm{d}t.$
Referring to the definition (2.5) of $\mathcal{E}$, (4.14) follows from (4.3). It then follows from the assumptions on $\mathbf{Q}$ that $\mathcal{E}(\mathbf{Q}^{n},\mathbf{Q}_{I,J,K,n}+\frac{r}{n}\mathbf{t})\geq 0$ for all $(I,J,K)\in T_{r}^{n}$, which by Proposition 3.7 guarantees that $\mathbf{Q}^{n}$ are the spectral quantiles of $n$-by-$n$ Hermitian matrices $A+B=C$. ∎
We are now equipped to prove our third main result, on self-characterisation.
See 2.3
Proof.
We first show that the second equivalence giving a criterion for membership in $\mathscr{H}$ follows from the first equivalence giving a criterion for membership in $\mathscr{H}^{\prime}$. Concretely, we prove that if $\mathcal{E}(\mathbf{Q},\mathbf{\tilde{Q}}+\mu\mathbf{t})\geq 0$ for all $\mathbf{Q}\in\mathscr{H}$, $\mathbf{\tilde{Q}}\in\mathscr{H}[0,1)$, and $0<\mu\leq\eta_{\mathbf{\tilde{Q}}}$, then in fact $\mathcal{E}(\mathbf{Q},\mathbf{\tilde{Q}}+\mu\mathbf{t})\geq 0$ for all $\mathbf{Q},\mathbf{\tilde{Q}}\in\mathscr{H}$ and $0\leq\mu\leq\eta_{\mathbf{\tilde{Q}}}$.
Note that if $\mathbf{Q}$ is a triple of bounded quantile functions and $\mathcal{E}(\mathbf{Q},\mathbf{\tilde{Q}}+\mu\mathbf{t})\geq 0$ for $0<\mu\leq\eta_{\mathbf{\tilde{Q}}}$, then sending $\mu\to 0$ we also obtain $\mathcal{E}(\mathbf{Q},\mathbf{\tilde{Q}})\geq 0$ by Lemma 4.3. Then, observing that the triples $\mathbf{\tilde{Q}}\in\mathscr{H}-\mathscr{H}[0,1)$ are precisely those for which $\eta_{\mathbf{\tilde{Q}}}=0$, it is enough to show that
(4.17) $\mathcal{E}(\mathbf{Q},\mathbf{\tilde{Q}})\geq 0\quad\forall\,\mathbf{Q}\in\mathscr{H},\ \mathbf{\tilde{Q}}\in\mathscr{H}[0,1)\quad\implies\quad\mathcal{E}(\mathbf{Q},\mathbf{\tilde{Q}})\geq 0\quad\forall\,\mathbf{Q},\mathbf{\tilde{Q}}\in\mathscr{H}.$
Let $\mathbf{Q},\mathbf{\tilde{Q}}\in\mathscr{H}$, and let $\bm{\pi}=(\pi_{1},\pi_{2},\pi_{3})$ be the triple of probability measures represented by $\frac{1}{2}\mathbf{Q}$. Then $\bm{\pi}\in\mathscr{H}$ by the dilation invariance property 2.3.2, and each $\pi_{i}$ is supported on [0, 1/2]012[0,\,1/2][ 0 , 1 / 2 ]. Let $\bm{\pi}^{\prime}=(\delta_{1/2},\,\delta_{1/2},\,\delta_{1})\in\mathscr{H}$ and write $\mathbf{Q}^{\prime}=(Q_{1}^{\prime},Q_{2}^{\prime},Q_{3}^{\prime})$ for the quantile functions of $\bm{\pi}^{\prime}$, that is,
$Q_{1}^{\prime}(t)$ $\equiv 1/2,$
$Q_{2}^{\prime}(t)$ $\equiv 1/2,$
$Q_{3}^{\prime}(t)$ $\equiv 1,\qquad\quad t\in[0,1].$
By the vertical convexity property 2.3.4, the average $\frac{1}{2}(\bm{\pi}+\bm{\pi}^{\prime})$ belongs to $\mathscr{H}$, with quantile functions $\mathbf{P}=(P_{1},P_{2},P_{3})$ given by
$P_{i}(t)=\begin{cases}\frac{1}{2}Q_{i}(2t),&t\in\big{[}0,\frac{1}{2}\big{)},\\ Q^{\prime}_{i}(2t-1),&t\in\big{[}\frac{1}{2},1\big{]}.\end{cases}$
Again by the dilation invariance property 2.3.2, $\frac{1}{2}\mathbf{\tilde{Q}}\in\mathscr{H}[0,1)$. From the definition (2.5) of $\mathcal{E}$, we find
$\mathcal{E}(\mathbf{Q},\mathbf{\tilde{Q}})=2\mathcal{E}\Big{(}\mathbf{P},\frac{1}{2}\mathbf{\tilde{Q}}\Big{)}$
where $\mathbf{P}\in\mathscr{H}$ and $\frac{1}{2}\mathbf{\tilde{Q}}\in\mathscr{H}[0,1)$. This proves (4.17), which reduces the second equivalence stated in the theorem to the first equivalence.
It remains to prove the criterion for membership in $\mathscr{H}^{\prime}$:
[…]
We first suppose that $\mathbf{Q}\in\mathscr{H}^{\prime}$ and show that the inequalities hold. By the definition of $\mathscr{H}^{\prime}$, we can choose a sequence $(\mathbf{Q}^{m})_{m\geq 1}$ of triples of quantile functions of empirical spectra of $m$-by-$m$ Hermitian matrices $A_{m}+B_{m}=C_{m}$, such that $\mathbf{Q}^{m}\to\mathbf{Q}$. Additionally, by Theorem 2.1 we can choose a sequence $(I_{n},J_{n},K_{n})_{n\geq 1}\in T_{p_{n}}^{r_{n}}$ such that $\mathbf{Q}_{I_{n},J_{n},K_{n},r_{n}}\to\mathbf{\tilde{Q}}$ and $p_{n}/r_{n}\to\mu$. Note that this implies
(4.18) $\mathbf{Q}_{I_{n},J_{n},K_{n},r_{n}}+\frac{p_{n}}{r_{n}}\mathbf{t}\ \to\ \mathbf{\tilde{Q}}+\mu\mathbf{t}.$
Now by Propositions 3.7 and 3.9, we have
$\mathcal{E}\Big{(}\mathbf{Q}^{m},\mathbf{Q}_{I_{n},J_{n},K_{n},r_{n}}+\frac{p_{n}}{r_{n}}\mathbf{t}\Big{)}\geq 0.$
Note that $\mathbf{Q}_{I_{n},J_{n},K_{n},r_{n}}+\frac{p_{n}}{r_{n}}\mathbf{t}$ represents a triple of probability measures with densities bounded above by $r_{n}/p_{n}$, and this ratio converges to $1/\mu$. In particular, there is some $n_{0}$ such that for all $n\geq n_{0}$, the density of each measure corresponding to a quantile function in the triple $\mathbf{Q}_{I_{n},J_{n},K_{n},r_{n}}+\frac{p_{n}}{r_{n}}\mathbf{t}$ is bounded above by $2/\mu$. Thus by (4.18) we may apply equation (4.1) of Lemma 4.1 to conclude that for each fixed $m$ we have
$\lim_{n\to\infty}\mathcal{E}\Big{(}\mathbf{Q}^{m},\mathbf{Q}_{I_{n},J_{n},K_{n},r_{n}}+\frac{p_{n}}{r_{n}}\mathbf{t}\Big{)}=\mathcal{E}(\mathbf{Q}^{m},\mathbf{\tilde{Q}}+\mu\mathbf{t}).$
In particular, since for each fixed $m$, every term in the sequence indexed by $n$ is nonnegative, the limit must also be nonnegative, that is,
$\mathcal{E}(\mathbf{Q}^{m},\mathbf{\tilde{Q}}+\mu\mathbf{t})\geq 0.$
We next note that $\mathbf{\tilde{Q}}+\mu\mathbf{t}$ represents a triple of measures with densities bounded above by $1/\mu$. Thus we may apply equation (4.2) of Lemma 4.1 to conclude that
$\lim_{m\to\infty}\mathcal{E}(\mathbf{Q}^{m},\mathbf{\tilde{Q}}+\mu\mathbf{t})=\mathcal{E}(\mathbf{Q},\mathbf{\tilde{Q}}+\mu\mathbf{t})\geq 0,$
where again the inequality holds because each term in the sequence is nonnegative. This completes the proof of the first direction of implication.
Now we prove the other direction. Namely, we show that if $\mathbf{Q}$ is a triple of quantile functions in $L^{1}([0,1])$ satisfying $\mathrm{tr}(\mathbf{Q})=0$ and with the property that for all $\mathbf{\tilde{Q}}\in\mathscr{H}[0,1)$ and $0<\mu\leq\eta_{\mathbf{\tilde{Q}}}$ we have $\mathcal{E}(\mathbf{Q},\mathbf{\tilde{Q}}+\mu\mathbf{t})\geq 0$, then $\mathbf{Q}\in\mathscr{H}^{\prime}$. In other words, we need to show that $\mathbf{Q}=(Q_{1}(t),Q_{2}(t),Q_{3}(t))$ arises as a limit of triples of quantile functions of empirical spectra of matrices $A_{n}+B_{n}=C_{n}$.
To this end, for each $n\geq 1$, $i=1,2,3$, and $t$ in an interval $\left[\frac{j-1}{n},\frac{j}{n}\right)$, we let $Q_{i}^{n}(t)$ be the average of $Q_{i}(t)$ on this interval:
$Q_{i}^{n}(t):=n\int_{(j-1)/n}^{j/n}Q_{i}(s)\,\mathrm{d}s\qquad\text{for }t\in\left[\frac{j-1}{n},\frac{j}{n}\right).$
Then $Q_{i}^{n}\to Q_{i}$ in $L^{1}$ by Lemma 4.6. Moreover by Proposition 2.8, for each $n\geq 1$, $\mathbf{Q}^{n}$ are the spectral quantile functions of a triple of Hermitian matrices $A_{n}+B_{n}=C_{n}$, and thus $\mathbf{Q}\in\mathscr{H}^{\prime}$. ∎
As a consequence of Theorem 2.3, we find that any triple of compactly supported probability measures satisfying the trace condition and all finite-$n$ Horn inequalities in fact satisfies all inequalities corresponding to points in the asymptotic Horn system.
Corollary 4.9.
Let $\mathbf{Q}$ be a triple of quantile functions of compactly supported measures with $\mathrm{tr}(\mathbf{Q})=0$. If $\mathbf{Q}$ satisfies the Horn inequalities, that is, if $\mathcal{E}(\mathbf{Q},\mathbf{Q}_{I,J,K,n}+\frac{r}{n}\mathbf{t})\geq 0$ for all $(I,J,K)$ in all $T^{n}_{r}$, then in fact $\mathcal{E}(\mathbf{Q},\mathbf{\tilde{Q}}+\mu\mathbf{t})\geq 0$ for all $\mathbf{\tilde{Q}}\in\mathscr{H}$ and $0\leq\mu\leq\eta_{\mathbf{\tilde{Q}}}$.
Proof.
If $\mathcal{E}(\mathbf{Q},\mathbf{Q}_{I,J,K,n}+\frac{r}{n}\mathbf{t})\geq 0$ for all $(I,J,K)$ in all $T^{n}_{r}$, the same holds for each $n$-average $\mathbf{Q}^{n}$, and thus $\mathbf{Q}^{n}$ represents the spectral quantiles of Hermitian matrices $A+B=C$. Since $\mathbf{Q}^{n}\to\mathbf{Q}$, we then have $\mathbf{Q}\in\mathscr{H}^{\prime}$.
The assumption that $\mathbf{Q}$ corresponds to a triple of compactly supported measures is equivalent to asssuming that each $Q_{i}$ is bounded. Let
$s$ $=\sup_{i=1,2,3;\ t\in[0,1]}\big{|}Q_{i}(t)\big{|}.$
We may assume that $s\neq 0$, since otherwise we would have $\mathbf{Q}=(0,0,0)\in\mathscr{H}$ and we would be done. The triple $\mathbf{Q}^{\prime}=\frac{1}{3s}(Q_{1}(t)+s,\,Q_{2}(t)+s,\,Q_{3}(t)+2s)$ belongs to $\mathscr{H}^{\prime}$ by the dilation and translation invariances 2.3.2, and each $Q^{\prime}_{i}$ takes values in $[0,1]$, so in fact $\mathbf{Q}^{\prime}\in\mathscr{H}$, and thus $\mathcal{E}(\mathbf{Q}^{\prime},\mathbf{\tilde{Q}}+\mu\mathbf{t})\geq 0$ for all $\mathbf{\tilde{Q}}\in\mathscr{H}$ and $0\leq\mu\leq\eta_{\mathbf{\tilde{Q}}}$ by Theorem 2.3. Since $\mathcal{E}(\mathbf{Q},\mathbf{\tilde{Q}}+\mu\mathbf{t})=3s\mathcal{E}(\mathbf{Q}^{\prime},\mathbf{\tilde{Q}}+\mu\mathbf{t})$, this proves the claim. ∎
Remark 4.10.
There are some similarities between the self-characterisation property and the notion of self-duality for cones in a finite-dimensional Euclidean space. We point this out in order to caution the reader, because the analogy only goes so far and may be misleading. These similarities become more evident if we keep track of the value of $\mu$ as an additional coordinate, because we then can state the self-characterisation property in such a way that each linear inequality on the quantile functions corresponds to a distinct point in a convex body. Define
[…]
$\mathcal{E}^{\bullet}\big{(}(\mathbf{Q},\mu),(\mathbf{\tilde{Q}},\tilde{\mu})\big{)}$ $:=\mathcal{E}(\mathbf{Q}+\mu\mathbf{t},\mathbf{\tilde{Q}}+\tilde{\mu}\mathbf{t})-\frac{1}{2}\mu\tilde{\mu}=\mathcal{E}(\mathbf{Q},\mathbf{\tilde{Q}}+\tilde{\mu}\mathbf{t}).$
Then for $\mathbf{Q}$ a triple of quantile functions taking values in $[0,1]$ with $\operatorname{tr}(\mathbf{Q})=0$, and for $0\leq\mu\leq\eta_{\mathbf{Q}}$, the self-characterisation property can be stated as
$(\mathbf{Q},\mu)\in\mathscr{H}^{\bullet}\iff\mathcal{E}^{\bullet}\big{(}(\mathbf{Q},\mu),(\mathbf{\tilde{Q}},\tilde{\mu})\big{)}\geq 0\quad\forall\,(\mathbf{\tilde{Q}},\tilde{\mu})\in\mathscr{H}^{\bullet}.$
The functional $\mathcal{E}^{\bullet}$ can be interpreted as a bilinear pairing via the two different linear structures described in Section 2.1: the linearity in the first argument comes from addition of quantile functions of probability measures on $\mathbb{R}$, while the linearity in the second argument comes from affine addition of finite signed measures on $[0,1]$. In that light, the above statement superficially resembles the definition of a self-dual cone. However, the analogy quickly breaks down. To begin with, although $\mathscr{H}^{\bullet}$ is convex, it is not a cone, due to the requirements that each $Q_{i}$ take values in $[0,1]$ and that $0\leq\mu\leq\eta_{\mathbf{Q}}$. Moreover, even though $\mathcal{E}^{\bullet}$ can be regarded as a bilinear pairing, it is a discontinuous and degenerate pairing between two different vector spaces, neither of which is contained in the topological dual of the other. To further complicate matters, the correspondence between quantile functions and probability measures is nonlinear, and $\mathcal{E}(\mathbf{Q},\mathbf{\tilde{Q}})$ is only defined when each $\tilde{Q}_{i}$ takes values in $[0,1]$, so the map
$(\mathbf{\tilde{Q}},\tilde{\mu})\mapsto\mathcal{E}^{\bullet}\big{(}\,\cdot\,,(\mathbf{\tilde{Q}},\tilde{\mu})\big{)}$
is not a linear transformation from the relevant vector space to its dual. The self-characterisation property of the asymptotic Horn system therefore does not correspond to a straightforward notion of “self-duality” as in finite-dimensional Euclidean spaces. Nevertheless, self-dual cones are self-characterising in the more general sense that we define in the following section; see Example 5.4 below.
5. Self-characterising sets and the question of uniqueness
In this section, we begin to develop a theory of sets that describe themselves in a fairly general sense. Concretely, we consider subsets $A\subset X$ where $X$ is an ambient set with a relation, such that relatedness with every element of $A$ is a necessary and sufficient condition for an element of $X$ to belong to $A$. We introduce and study both a weak and a strong form of self-characterisation.
The motivation for this theory is the question: how close does the self-characterisation property (Theorem 2.3) come to uniquely determining the asymptotic Horn system $\mathscr{H}$? It is too much to hope that Theorem 2.3 alone could be used to define $\mathscr{H}$, but it can form part of a definition when combined with some additional information, and we prove a modest uniqueness result in this direction. We also prove that weak self-characterisation by itself does not allow any reduction of the relevant Horn inequalities, in the sense that if some subset of Horn inequalities implies all of the others under the assumption that the full set of inequalities is weakly self-characterising, then that subset implies the others even without the assumption of self-characterisation. We conjecture that a similar statement is true for strong self-characterisation. It remains possible that weak or strong self-characterisation plus further assumptions of a geometric or topological nature (such as the self-averaging property and the fact that $(n,r)$-lattice points are dense in $\mathscr{H}$) might be sufficient to uniquely determine the asymptotic Horn system.
5.1. Generalities
We turn now to the more abstract setting of a set endowed with a relation.
Definition 5.1.
Let $X$ be a set, and let $L\subset X\times X$ be a relation on $X$. We write $xLy$, or say $x$ likes $y$, if $(x,y)\in L$. Otherwise we say $x$ dislikes $y$. We say that $x$ is self-liking if $xLx$, otherwise $x$ is self-disliking.
We say that a subset $A\subset X$ is friendly (under $L$) if
(5.1) $\forall x,y\in A\ \ xLy.$
We say that $A$ is weakly packed (under $L$) if
(5.2) $\forall x\in X\ \ \big{(}(\forall y\in A\cup\{x\}\ \ xLy\ \&\ yLx)\implies x\in A\big{)}.$
Equivalently, $A$ is weakly packed if every $x$ not in $A$ either dislikes or is disliked by some $y$ in $A$, or is self-disliking.
We say that $A$ is strongly packed (under $L$) if
(5.3) $\forall x\in X\ \ \big{(}(\forall y\in A\ \ xLy)\implies x\in A\big{)}.$
Equivalently, $A$ is strongly packed if every $x$ not in $A$ dislikes some $y$ in $A$. Clearly, a set that is strongly packed is also weakly packed, but if $L$ is not symmetric and reflexive then the converse is not necessarily true.
We say that $A$ is weakly self-characterising (under $L$) if $A$ is both friendly and weakly packed under $L$. Equivalently,
(5.4) $A=\{x\in X\ |\ \forall y\in A\cup\{x\}\ \ xLy\ \&\ yLx\}.$
Analogously, we say that $A$ is strongly self-characterising (under $L$) if $A$ is both friendly and strongly packed under $L$, or equivalently,
(5.5) $A=\{x\in X\ |\ \forall y\in A\ \ xLy\}.$
The name self-characterising refers to the fact that relatedness with all elements of a self-characterising set (or mutual relatedness together with self-liking) provides a necessary and sufficient condition for membership.
[…]
Figure 1. A relation on the set $X=\{x_{1},\ldots,x_{5}\}$. We draw a green arrow from $x_{i}$ to $x_{j}$ if $x_{i}$ likes $x_{j}$. The subset $A=\{x_{1},x_{2},x_{3}\}$ is strongly self-characterising, as all elements of $A$ like each other and themselves, and every element not in $A$ dislikes an element of $A$.
If $A,B\subseteq X$ are two subsets, then we say that $A$ is weakly or strongly packed (resp. self-characterising) under $L$ relative to $B$ if $A\cap B$ is weakly or strongly packed (resp. self-characterising) under the restriction of $L$ to $B$. There is no need for a notion of friendliness relative to $B$, since the property of friendliness depends only on $A$ and not on a choice of ambient set.
We will sometimes suppress explicit mention of the relation $L$ when it is clear from context.
We stress that self-characterisation can only be an informative property once we have fixed a particular relation. Otherwise it tells us nothing whatsoever about the set $A$. Indeed, any $A\subset X$ is strongly self-characterising under the relation $L=A\times A\subset X\times X$. On the other hand, if the set $X$ and the relation $L$ have some additional structure, then self-characterisation may automatically guarantee other properties, as illustrated in the following easy example.
Proposition 5.2.
Let $X$ be a topological space, let $f:X\times X\to\mathbb{R}$ be a continuous function, and consider the relation $L$ defined by $xLy\iff f(x,y)\geq 0$. If $A\subset X$ is weakly self-characterising, then $A$ is closed.
Proof.
Since $A$ is weakly self-characterising, we have
$A$ =$=\{x\in X\ |\ \forall y\in A\ \ f(x,y),f(y,x),f(x,x)\geq 0\}$
$=\bigcap_{y\in A}\{x\in X\ |\ f(x,y)\geq 0\}\cap\bigcap_{y\in A}\{x\in X\ |\ f(y,x)\geq 0\}\cap\{x\in X\ |\ f(x,x)\geq 0\}.$
Since $f$ is continuous, the functions $x\mapsto f(x,y)$, $x\mapsto f(y,x)$, and $x\mapsto f(x,x)$ are continuous as well, and thus their upper contour sets are closed. Therefore $A$ is an intersection of closed sets and must itself be closed. ∎
Moreover, for certain types of relations, self-characterisation is equivalent to more familiar properties.
Example 5.3.
If $L$ is an equivalence relation, then weakly self-characterising subsets are also strongly self-characterising and are precisely the equivalence classes of $L$.
Example 5.4.
If $X$ is a Euclidean space and $L$ is the relation defined by
$xLy\quad\iff\quad\langle x,y\rangle\geq 0,$
then the self-characterising subsets are the self-dual cones in $X$. Again there is no distinction between weak and strong self-characterisation in this example, since the inner product is symmetric and positive definite, and therefore $L$ is symmetric and reflexive.
Friendly, packed, and self-characterising sets satisfy some convenient properties with regard to containment of subsets.
Proposition 5.5.
Let $X$ be a set endowed with a relation $L$, and let $A\subsetneq B\subset X$ be subsets.
  1. (1)
    If $B$ is friendly, then $A$ is also friendly and is not weakly (or strongly) packed.
  2. (2)
    If $A$ is weakly (resp. strongly) packed, then $B$ is also weakly (resp. strongly) packed and is not friendly.
  3. (3)
    The weakly self-characterising subsets of $X$ are precisely the sets that are friendly and are not contained in another friendly set (i.e. the sets that are maximal under containment among friendly sets), or equivalently the sets that are weakly packed and do not strictly contain another weakly packed set (i.e. the sets that are minimal under containment among weakly packed sets).
Proof.
To prove (1), suppose that $B$ is friendly. For any $x,y\in A$, we also have $x,y\in B$, and therefore $xLy$, so that $A$ is friendly as well. On the other hand, since $A\neq B$, there must be some element $x\in B-A$. We thus have $xLy$ and $yLx$ for all $y\in A$, but $x\not\in A$, so $A$ is not weakly packed.
To prove (2), suppose that $A$ is weakly (resp. strongly) packed. For any $x\in X$, if $xLy$ (resp. $xLy$, $yLx$ and $xLx$) for all $y\in B$, then in particular $xLy$ (resp. $xLy$ and $yLx$) for all $y\in A$. Since $A$ is weakly (resp. strongly) packed, we must then have $x\in A$, and thus also $x\in B$, and therefore $B$ is weakly (resp. strongly) packed as well. On the other hand, for $x\in B-A$, either $x$ must be self-disliking or there must be some $y\in A$ that either dislikes or is disliked by $x$, since otherwise $A$ would not be weakly packed. Therefore $B$ is not friendly.
To prove (3), suppose that $A$ is a weakly self-characterising set. Then $A$ is both weakly packed and friendly, and thus property (1) guarantees that $A$ cannot strictly contain another weakly packed set, while (2) guarantees that $A$ cannot be strictly contained in another friendly set. Next suppose that $A$ is a friendly set that is maximal under containment, and suppose for the sake of contradiction that $A$ is not weakly self-characterising (i.e. not weakly packed). Then there must exist some $x\in X-A$ that likes itself and also likes and is liked by all elements of $A$. But then $A\cup\{x\}$ would be a friendly set that strictly contains $A$, contradicting the assumption that $A$ is maximal. Finally, suppose that $A$ is a weakly packed set that is minimal under containment, and suppose for the sake of contradiction that $A$ is not friendly. Then there must be some $x,y\in A$, possibly not distinct, such that $x$ dislikes $y$. But then $A-\{x\}$ would be a weakly packed strict subset of $A$, contradicting the assumption that $A$ is minimal. ∎
We now turn to questions of existence and uniqueness of self-characterising sets. In the proof of the following corollary, in the setting where $X$ is infinite we assume the axiom of choice.
Corollary 5.6.
Let $X$ be a set, and $L$ a relation on $X$. Then $X$ contains at least one weakly self-characterising set.
More generally, if $E$ is any friendly subset of $X$, there exists a weakly self-characterising set xxcontaining $E$.
Proof.
First observe that the empty subset is friendly under any relation on any set, so the friendly subsets of $X$, with the partial order of containment, form a non-empty poset. If $X$ is finite, then this poset must contain a maximal element, which by Proposition 5.5 is a self-characterising subset. The same argument applies if we restrict our attention to the poset of friendly subsets containing a given friendly subset $E$.
If $X$ is infinite, then the same conclusion follows by invoking Zorn’s lemma. If $A_{1}\subseteq A_{2}\subseteq\ldots\subseteq X$ is an ascending chain of friendly subsets, then the union $\bigcup_{n=1}^{\infty}A_{n}$ is also friendly and is an upper bound for the chain. By Zorn’s lemma, there is thus a friendly subset of $X$ that is maximal under containment, and by Proposition 5.5 this subset is self-characterising. ∎
We can also ask when there exists a unique weakly self-characterising set containing a given friendly set. The proof of the following proposition again requires the axiom of choice when the set $X$ is infinite.
Proposition 5.7.
Let $L$ be a relation on a set $X$, let $E$ be a friendly subset of $X$, and let $V$ be a weakly self-characterising set containing $E$. The following are equivalent:
  1. (1)
    $V$ is the unique weakly self-characterising set containing $E$.
  2. (2)
    For $x\in X$ self-liking,
    (5.6) $\big{(}\forall y\in E\ \ xLy~{}\&~{}yLx\big{)}\iff\big{(}\forall y\in V\ \ xLy~{}\&~{}yLx\big{)}.$
Proof.
First we prove that (2) implies (1) by considering the contrapositive: we show that if (1) does not hold, then (2) does not hold. In this direction, suppose there are distinct weakly self-characterising sets $V$ and $W$, both of which contain $E$. By Proposition 3.2, neither $V$ nor $W$ contains the other. In particular, there exists $x\in W-V$. Since $W$ is friendly, $x$ is self-liking and both likes and is liked by all elements of $W$, and in particular, $x$ likes and is liked by all elements of $E$. However, $x$ cannot both like and be liked by all elements of $V$, since otherwise $x$ would have to belong to $V$ because $V$ is weakly packed. Thus (5.6) does not hold.
We now show (1) implies (2). Suppose $V$ is the unique weakly self-characterising set containing $E$, and suppose $x$ is some self-liking element of $X$ satisfying $xLy$ and $yLx$ for all $y\in E$. If $x$ itself is in $V$, then $x$ likes and is liked by all elements of $V$, and we are done.
We now show that if $x$ is not in $V$, then we contradict (1). Indeed, if $x$ is not in $V$, then $x$ is a self-liking element that likes and is liked by all elements of $E$. Thus $E\cup\{x\}$ is friendly, and by Corollary 5.6, there exists some weakly self-characterising set $W$ containing $E\cup\{x\}$. Note that $x\in W-V$, thus $W$ and $V$ are distinct wealky self-characterising sets containing $E$, a contradiction. ∎
Taking $E$ to be the empty set in Proposition 5.7, the following is immediate:
Corollary 5.8.
Let $L$ be a relation on a set $X$, and suppose that $V\subseteq X$ is the unique weakly self-characterising subset under $L$. Then $V=\{x\in X\ |\ xLx\}.$
The existence and uniqueness of strongly self-characterising sets is a different matter altogether. An analogous statement to Corollary 5.6 does not hold for strongly self-characterising sets: it is not true in general that every friendly set is contained in a strongly self-characterising set, and indeed strongly self-characterising sets may not exist at all, as shown in the example in Figure 2.
$x$$y$
Figure 2. A relation on the set $X=\{x,y\}$ for which there are no strongly self-characterising sets. The friendly sets are the empty set and the singleton $\{x\}$, while the strongly packed sets are the singleton $\{y\}$ and the full set $\{x,y\}$.
5.2. The Horn inequalities and the question of uniqueness
We now apply the preceding discussion to the Horn inequalities.
[…]
$\mathbf{Q}\Vdash\mathbf{\tilde{Q}}\quad\iff\quad\mathcal{E}(\mathbf{Q},\mathbf{\tilde{Q}}+\mu\mathbf{t})\geq 0\quad\forall\,\mu\in[0,\eta_{\mathbf{\tilde{Q}}}].$
Define
(5.7) $\begin{split}\mathcal{U}_{n}&=\bigcup_{m=2}^{n}\bigcup_{r=1}^{m-1}\Big{\{}\,\mathbf{Q}_{I,J,K,m}\ \ \Big{|}\ \ (I,J,K)\in U^{m}_{r}\,\Big{\}},\\ \mathcal{U}&=\bigcup_{n=2}^{\infty}\mathcal{U}_{n},\end{split}$
with $U^{n}_{r}$ defined by (3.2) and $\mathbf{Q}_{I,J,K,m}$ defined by (3.6) and (3.7). Then we have the following modest uniqueness result.
Theorem 5.9.
[…]
Proof.
[…]
Observe that
$\mathscr{H}\cap\mathcal{U}_{n}=\Big{\{}\mathbf{Q}_{I,J,K,m}\ \ \Big{|}\ \ (I,J,K)\in T^{m}_{r}\text{ for some $1\leq r<m\leq n$}\Big{\}},$
and that for any such triple $(I,J,K)$, $\mathbf{Q}_{I,J,K,m}+\frac{r}{m}\mathbf{t}\in\mathscr{H}[0,1)$. Thus for any $\mathbf{Q}\in\mathcal{U}_{n}$, by Theorem 2.3 and Propositions 3.7 and 3.9,
$\mathbf{Q}\Vdash\mathbf{\tilde{Q}}\quad\forall\,\mathbf{\tilde{Q}}\in\mathscr{H}\cap\mathcal{U}_{n}\quad\iff\quad\mathcal{E}\Big{(}\mathbf{Q},\mathbf{Q}_{I,J,K,m}+\frac{r}{m}\mathbf{t}\Big{)}\geq 0\quad\forall\,(I,$ $J,K)\in T^{m}_{r},$
$1\leq r<m\leq n.$
The fact that $\mathscr{H}$ is weakly self-characterising relative to $\mathcal{U}_{n}$ for all $n\geq 2$ then follows from Lemma 3.5 and Proposition 3.9.
To see that $\mathscr{H}$ is the unique weakly self-characterising subset of $\mathcal{M}$ with this property, let $S\subset\mathcal{M}$ be any weakly self-characterising set that is also weakly self-characterising relative to $\mathcal{U}_{n}$ for all $n\geq 2$. We will show that $S=\mathscr{H}$.
First observe that it suffices to prove that $S\cap\mathcal{U}_{n}=\mathscr{H}\cap\mathcal{U}_{n}$. We then have
$\bigcup_{n=2}^{\infty}\mathscr{H}\cap\mathcal{U}_{n}\subseteq S,$
and since $S$ is friendly, every $\mathbf{Q}\in S$ must satisfy
[…]
Calculating as in Proposition 2.8, we find that the same must hold for each averaged triple $\mathbf{Q}^{n}$, and thus $\mathbf{Q}^{n}$ represents the empirical spectral measures of an $n$-by-$n$ Hermitian triple $A_{n}+B_{n}=C_{n}$ with all eigenvalues in $[0,1]$. Since $\mathbf{Q}^{n}\to\mathbf{Q}$, we find that $\mathbf{Q}$ is a limit of spectra of Hermitian triples, and thus $\mathbf{Q}\in\mathscr{H}$ by definition, implying $S\subseteq\mathscr{H}$. Since one weakly self-characterising set cannot strictly contain another by Proposition 5.5, we can then conclude that $S=\mathscr{H}$.
Therefore it only remains to prove that $S\cap\mathcal{U}_{n}=\mathscr{H}\cap\mathcal{U}_{n}$, which we do by induction. The base case is $n=2$. From (3.2) we have
$U^{2}_{1}=\Big{\{}\big{(}\{1\},\{1\},\{1\}\big{)},\ \big{(}\{1\},\{2\},\{2\}\big{)},\ \big{(}\{2\},\{1\},\{2\}\big{)}\Big{\}},$
corresponding to the Weyl inequalities for 2-by-2 matrices. A direct calculation shows that the set
$\mathcal{U}_{2}=\big{\{}\mathbf{Q}_{I,J,K,2}\ \big{|}\ (I,J,K)\in U^{2}_{1}\big{\}}=\big{\{}\mathbf{Q}_{I,J,K,2}\ \big{|}\ (I,J,K)\in T^{2}_{1}\big{\}}$
[…]
Suppose now that $S\cap\mathcal{U}_{k}=\mathscr{H}\cap\mathcal{U}_{k}$ for some $k\geq 2$. We will show that $S\cap\mathcal{U}_{k+1}=\mathscr{H}\cap\mathcal{U}_{k+1}$. Again by by Proposition 5.5, it suffices to show that $S\cap\mathcal{U}_{k+1}\subseteq\mathscr{H}\cap\mathcal{U}_{k+1}$. Let $\mathbf{Q}\in S\cap\mathcal{U}_{k+1}$. Then $\mathbf{Q}\Vdash\mathbf{\tilde{Q}}$ for all $\mathbf{\tilde{Q}}\in\mathscr{H}\cap\mathcal{U}_{k}$, since $\mathscr{H}\cap\mathcal{U}_{k}=S\cap\mathcal{U}_{k}\subset S\cap\mathcal{U}_{k+1}$, and $S\cap\mathcal{U}_{k+1}$ is friendly. By Lemma 3.5 and Proposition 3.9, this means that $\mathbf{Q}=\mathbf{Q}_{I,J,K,m}$ for some $(I,J,K)\in T^{m}_{r}$ with $11\leq r<m\leq k+$ 1, so that $\mathbf{Q}\in\mathscr{H}\cap\mathcal{U}_{k+1}$, and thus $S\cap\mathcal{U}_{k+1}\subseteq\mathscr{H}\cap\mathcal{U}_{k+1}$ as desired. ∎
The proof of Theorem 5.9 relies on the inductive structure of the sets $T^{n}_{r}$, but one could hope to find a condition that uniquely determines $\mathscr{H}$ as a self-characterising set without exploiting this structure. For instance, one could look for subsets $E\subset\mathscr{H}$ such that $\mathscr{H}$ is the unique set that is (weakly or strongly) self-characterising and contains $E$. Such subsets clearly exist, since the property holds if we take $E=\mathscr{H}\cap\mathcal{U}$ to be the set of all triples representing Horn inequalities. But this trivial choice cannot be minimal, since the Horn inequalities themselves are known to be redundant for $n\geq 5$ (see [6, 18]), and the property must also hold if we take $E$ to be a maximal non-redundant subset of $\mathscr{H}\cap\mathcal{U}$. In fact, assuming the axiom of choice, Proposition 5.7 above immediately implies that if we are interested in describing $\mathscr{H}$ as the unique weakly self-characterising set containing some given set $E$, then finding such a set $E$ that is minimal under containment is equivalent to finding a maximal non-redundant subset of Horn inequalities:
Corollary 5.10.
Let $E\subset\mathscr{H}$ be a subset that uniquely determines $\mathscr{H}$ as a weakly self-characterising set, in the sense that if $S\subset\mathcal{M}$ is weakly self-characterising and $E\subset S$, then $S=\mathscr{H}$. Then for all $\mathbf{Q}\in\mathcal{M}$,
$\mathbf{Q}\Vdash\mathbf{\tilde{Q}}\quad\forall\,\mathbf{\tilde{Q}}\in E\quad\iff\quad\mathbf{Q}\Vdash\mathbf{\tilde{Q}}\quad\forall\,\mathbf{\tilde{Q}}\in\mathscr{H}.$
We conjecture that a similar statement holds if one considers strong rather than weak self-characterisation.
6. Directions for further work
Even with the above results on self-characterisation, dense subsets, and other properties of the set $\mathscr{H}$, the geometry of the asymptotic Horn system remains fairly mysterious, and we suspect the reader may agree that the preceding analysis has raised as many questions as it has answered. To close, we point out some directions for future research that would be natural sequels to the investigation in this paper.
First, some of the results in this article can likely be improved. For example, it is not immediately apparent whether Theorem 4.7 on the redundancy in the Horn inequalities in the $n\to\infty$ limit is sharp. It is natural to ask whether a converse statement holds: given sequences $r_{k}<n_{k}$ such that $(r_{k}/n_{k})_{k\geq 1}$ is not dense in $(0,1)$, can one construct Horn inequalities in some $T^{n}_{r}$ that are not implied by the inequalities in $(T^{n_{k}}_{r_{k}})_{k\geq 1}$? And we have not touched the more detailed question of which further inequalities in each $T^{n_{k}}_{r_{k}}$ could additionally be discarded without changing the set of solutions.
[…]
Another possible direction of inquiry is to relate the present work to probabilistic versions of Horn’s problem in the $n\to\infty$ limit. Recent work in random matrix theory has studied random matrix ensembles that can be regarded as quantitative versions of Horn’s problem, in which one would like to know not only whether a given triple $(\alpha,\beta,\gamma)\in\mathbb{R}^{3n}$ can occur as the spectra of Hermitian matrices $A+B=C$, but also, in a certain sense, how many such matrices there are [13, 12, 11]. It is natural to wonder whether one could formulate a large-$n$ version of the randomised Horn’s problem in which $\mathscr{H}$ or the asymptotic Horn bodies $\mathscr{H}_{\mu,\nu}$ defined in (2.2) would play the role of the finite-dimensional Horn polytopes. This question is also related to recent work of Narayanan, Sheffield and Tao on asymptotics of random hives [28, 29], since the probability measure studied in the randomised Horn’s problem is the pushforward by a linear projection of the uniform probability measure on a hive polytope.
Finally, Horn’s problem and its probabilistic version can both be viewed from a much more general perspective; namely, they are special cases of the problems of determining the moment polytope and Duistermaat–Heckman measure of a symplectic manifold equipped with a Hamiltonian group action [15, 22]. A number of other well-known random matrix ensembles can be viewed in the same light, including the randomised Schur’s problem, the randomised quantum marginals problem, uniform random antisymmetric matrices with deterministic eigenvalues, and the Hermitian orbital corners process [11, 12, 14, 26, 7]. Asymptotic questions about all of the above models can be interpreted as problems in high-dimensional symplectic geometry, and one could try to ask and answer interesting questions about moment polytopes and Duistermaat–Heckman measures arising from actions of high-dimensional Lie groups in a more general setting than the particular case of Horn’s problem.
Acknowledgements
The work of C.M. is supported by the National Science Foundation under grant number DMS-2103170 and by a grant from the Simons Foundation.